### v3

# Tail Index Estimation for Discrete Heavy-Tailed Distributions with Application to Statistical Inference for Regular Markov Chains

Abstract
It is the purpose of this paper to investigate the issue of estimating the regularity index $\beta>0$ of a discrete heavy-tailed r.v. $S$, i.e. a r.v. $S$ valued in $\mathbb{N}^{*}$ such that $\mathbb{P}(S>n)=L(n)\cdot n^{-\beta}$ for all $n\geq 1$, where $L:\mathbb{R}^{*}_{+}\to\mathbb{R}_{+}$ is a slowly varying function. Such discrete probability laws, referred to as generalized Zipf’s laws sometimes, are commonly used to model rank-size distributions after a preliminary range segmentation in a wide variety of areas such as e.g. quantitative linguistics, social sciences or information theory. As a first go, we consider the situation where inference is based on independent copies $S_{1},\;\ldots,\;S_{n}$ of the generic variable $S$. Just like the popular Hill estimator in the continuous heavy-tail situation, the estimator $\widehat{\beta}$ we propose can be derived by means of a suitable reformulation of the regularly varying condition, replacing $S$’s survivor function by its empirical counterpart. Under mild assumptions, a non-asymptotic bound for the deviation between $\widehat{\beta}$ and $\beta$ is established, as well as limit results (consistency and asymptotic normality). Beyond the i.i.d. case, the inference method proposed is extended to the estimation of the regularity index of a regenerative $\beta$-null recurrent Markov chain. Since the parameter $\beta$ can be then viewed as the tail index of the (regularly varying) distribution of the return time of the chain $X$ to any (pseudo-) regenerative set, in this case, the estimator is constructed from the successive regeneration times. Because the durations between consecutive regeneration times are asymptotically independent, we can prove that the consistency of the estimator promoted is preserved. In addition to the theoretical analysis carried out, simulation results provide empirical evidence of the relevance of the inference technique proposed.
keywords:
generalized discrete Pareto distribution, nonparametric estimation, null recurrent Markov chain, regularity index, Zipf’s law
pacs:
MSC Classification]60K35
1 Introduction
This article is devoted to the study of the problem of estimating the regularity index $\beta>0$ of a generalized discrete Pareto distribution, namely the probability distribution of a random variable $S$ defined on a probability space $(\Omega,\;\mathcal{F},\;\mathbb{P})$, taking its values in $\mathbb{N}^{*}$ and such that:
$\mathbb{P}\left({S>n}\right)={n^{-\beta}}L\left(n\right)\text{ for all }n\geq 1,$ (1)
where $L:\mathbb{R}_{+}\to\mathbb{R}$ is a slowly varying function, i.e. such that $L(\lambda z)/L(z)\rightarrow+1$ as $z\rightarrow+\infty$ for any $\lambda>0$, see [7]. Such discrete power law probability distributions, also referred to as generalized Zipf’s laws sometimes, are often used to model the distribution of discrete data exhibiting a specific rank-frequency relationship, namely when the logarithm of the frequency and that of the rank order are nearly proportional. Such a phenomenon has been empirically observed in many ranking systems: in quantitative linguistics (i.e. when analysing word frequency law in natural language, see e.g. [30]) in the first place, as well as in a very wide variety of situations, too numerous to be exhaustively listed here. One may refer to [37], [26] or [40] among many others. In this paper, we first consider the issue of estimating the parameter $\beta$ involved in (1) (supposedly unknown, like the function $L$) in the classic (asymptotic) i.i.d. statistical setting, i.e. based on an increasing number $n\geq 1$ of independent copies $S_{1},\;\ldots,\;S_{n}$ of the generic r.v. $S$. Statistical inference for discrete heavy-tailed distributions has not received much attention in the literature and still poses methodological problems that this article seeks to resolve. Most of the very few dedicated methods documented either deal with very specific cases as in e.g. [15], [28] and [11] or else consist in empirically applying techniques originally designed for continuous heavy-tailed distributions to discrete data after first adding independent uniform noise [39, 21, 22].
The vast majority of the regular variation index estimators proposed in the literature, Hill or Pickand estimators in particular [19, 35], are based on order statistics, which causes obvious difficulties in the discrete case because of the possible occurrence of many ties. Indeed, in that case, many spacings are equal to 00 which causes these estimators to behave erroneously. This is stressed in [28] (see their figures 1 and 2). More importantly, [28] constructs an example of a family of discrete value heavy-tailed distributions that do not satisfy any second-order condition so that the asymptotic normality of the Hill estimator cannot be proved directly. In contrast, the estimator under study here is based on the analysis of the probability of exponentially separated tail events. It simply rests on the fact that, as can be immediately deduced from (1), we have $\ln(p_{k})-\ln(p_{k+1})=\beta+\ln(L(e^{k})/L(e^{k+1}))$, where $\ln(x)$ denotes the natural logarithm of any real number $x>0$ and $p_{l}=\mathbb{P}(S>e^{l})$ for all $l\in\mathbb{N}$, and that $L(e^{k+1})/L(e^{k})$ is expected to be very close to $1$ for $k\in\mathbb{N}$ chosen sufficiently large. A natural (plug-in) inference technique can be then devised by replacing the tail probabilities $p_{l}$ with their empirical versions $\widehat{p}^{(n)}_{l}=(1/n)\sum_{i=1}^{n}\mathbb{I}\{S_{i}>e^{l}\}$ for $l\in\mathbb{N}$, where $\mathbb{I}\{\mathcal{A}\}$ means the indicator function of any event $\mathcal{A}$. This yields the estimator
$\widehat{{\beta}}_{n}\left(k\right)=\ln\left({\widehat{p}_{k}^{(n)}}\right)-\ln\left(\widehat{p}_{k+1}^{(n)}\right),$ (2)
provided that $\widehat{p}_{k+1}^{(n)}>0$ (as shall be seen, this occurs with large probability if $n$ is sufficiently large). By convention, we set $\widehat{{\beta}}_{n}(k)=0$ when $\widehat{p}_{k+1}^{(n)}=0$. We point out that it has exactly the same form as that proposed and analysed in [10] in a different context, that of (continuous) approximately Pareto distributions [^1] The distribution of a real-valued r.v. $X$ is said to be approximately Pareto with tail index $\beta>0$ if its survivor function is of the form: $\forall x>0$, $\mathbb{P}(X>x)=L(x)x^{-\beta}$, where $L$ is asymptotically constant at infinity, i.e. there exists $C\in(0,\;\infty)$ s.t. $L(x)\to C$ as $x\rightarrow+\infty$. namely. In the discrete generalized Pareto framework, we prove that for an appropriate choice of the hyper-parameter $k=k_{n}$ (typically chosen of order $\ln(n)$), the estimator (2) is strongly consistent and asymptotically normal as $n\rightarrow+\infty$. We also show that non-asymptotic upper confidence bounds for the absolute deviations between $\widehat{{\beta}}_{n}\left(k\right)$ and $\beta$ are also established here.
Although estimation of the parameter $\beta$ in (1) in the discrete i.i.d. setting is an important issue in itself, the present paper also finds its motivation in the problem of recovering statistically the regularity index of a regenerative regular $\beta$-null-recurrent Markov chain $X=(X_{n})_{n\in\mathbb{N}}$, based on the observation of a finite sample path $X_{1},\;\ldots,\;X_{n}$ with $n\geq 1$. As explained in [8, 9], for regular Markov chains, the regularity index $\beta$ controls the (sublinear) rate at which the number of visits to any given Harris set increases with observation time $n$, no matter the initial distribution. These Markov chains are strongly non-stationary and are not mixing in any sense. Typical examples of these chains are random walks and their variations, for instance Bessel random walks (see [2]). Such processes appear also naturally in many applications: for population dynamics or stochastic growth models in demography, see for instance [1] and the references therein; for branching processes, Galton-Watson processes with immigration (or population dependent) also exhibit a fat tail behaviour of the time returns of the chain, see [34, 41, 23].
In the regenerative case (i.e. when the chain $X$ possesses an accessible atom, a Harris set on which the transition probability is constant), the distribution of the regenerative time, the return time to the atom, is a discrete generalized Pareto (1) and the parameter $\beta$ is its tail index. Due to the non-standard behaviour of traditional estimators in this context, statistical inference for null-recurrent Markov chains is very poorly documented in the literature [25, 24, 27, 17] and, to the best of our knowledge, estimation of the key quantity $\beta$ has not received much attention. It is also the goal of this article to extend the use of the estimator (2) to the case where the $S_{i}$’s are the successive durations between the consecutive regeneration times up to time $n$. The main difficulty naturally arises from the fact that the number $1+N_{n}\geq 0$ of regeneration times (and thus the number of durations) is now random, and it is not independent of the variables $S_{1},\ldots,S_{N_{n}}$ (in particular, $S_{1}+\cdots+S_{N_{n}}\leq n$ by construction). We show that the estimator remains strongly consistent in this scenario and we provide non-asymptotic upper confidence bounds for the absolute deviations between $\widehat{{\beta}}_{N_{n}}\left(k\right)$ and $\beta$. For illustration purposes, numerical experiments have been carried out, providing empirical evidence of the relevance of the estimation method promoted. Extension to the general case of (pseudo-regenerative) null-recurrent chains is also discussed, the difficulties inherent in applying the methodology originally proposed in [6] in the positive recurrent case to mimic regenerative Nummelin extensions [33] being explained at length.
The paper is organized as follows. A thorough analysis of the behaviour of the estimator (2) in the i.i.d. case, illustrated by numerical experiments, is first carried out in section 2. Some of the results thus established (consistency and non-asymptotic bounds) are next extended in section 3 to the regenerative regular Markovian setup, when the estimator is computed based on a single finite-length trajectory of the atomic chain. Experimental results are also displayed and the main barrier to the extension of the methodology promoted to general (i.e. pseudo-regenerative) regular null-recurrent chains is also discussed therein. Technical proofs are deferred to the Appendix section.
2 Tail Index Estimation - The Discrete Heavy-Tailed i.i.d. Case
Throughout this section, $S_{1},\ldots,S_{n}$ are independent copies of a generic discrete generalized Pareto r.v. $S$, i.e. a r.v. $S$ with survivor function of type (1), where the parameter $\beta>0$ and the slowly varying function $L$ are supposedly unknown. As a first go, we start to investigate the (asymptotic) behaviour of the estimator (2) in this basic general framework and next develop the analysis in particular situations, i.e. when the function $L$ has a specific form.
2.1 Main Results - Confidence Bounds and Limit Theorems
As explained in the Introduction section, the estimator (2) can be viewed as an empirical counterpart of the quantity
$\beta(k)=\ln(p_{k})-\ln(p_{k+1})=\beta+\ln\left(\frac{L(e^{k})}{L(e^{k+1})}\right),$ (3)
see (1), which tends to $\beta$ as $k\rightarrow\infty$ by virtue of the slow variation property of $L$. As previously emphasized, unless the function $L$ is supposed to be asymptotically constant (i.e. there exists $C>0$ s.t. $L(x)\rightarrow C$ as $x\rightarrow+\infty$), the discrete generalized Pareto model (1) is not a discrete version of the (continuous) approximately $\beta$-Pareto model considered in [10] and, consequently, the validity framework established therein does not apply directly here. The proposition below provides an upper confidence bound for the absolute deviations between (2) and $\beta$ (respectively, between (2) and $\beta(k)$).
Proposition 2.1.
[…]
$\left|\widehat{\beta}_{n}(k)-\beta\right|\leq 6\sqrt{\frac{u_{n}(\delta)}{p_{k+1}}}+\left|\ln\left(\frac{L(e^{k})}{L(e^{k+1})}\right)\right|.$ (4)
Refer to the Appendix section for the technical proof. The bound (4) reveals some sort of “’bias-variance” trade-off, ruled by the hyperparameter $k>0$. The second term on the right-hand side can be viewed as the bias of the inference method, insofar as the estimator (2) can be seen as an empirical version of the approximation (3). It decays to 00 as $k$ increases towards infinity, while the first term, whose presence is due to the random nature of the estimator, tends to +$\infty$+ ∞. We point out that second-order slow variation conditions [16] are required to bound the (vanishing) bias term in (4), as shall be explained in subsection 2.2. The following result reveals that for an appropriate choice of $k=k_{n}$, the estimator (2) is strongly consistent.
Theorem 2.1 (Strong consistency).
Suppose that, as $n\rightarrow+\infty$, we have $k_{n}\to+\infty$ and $(\ln n)\exp(k_{n}\beta)/n=o(L(\exp(k_{n}))$. Then, we have:
$\widehat{\beta}_{n}(k_{n})\to\beta\text{ almost surely, as }n\rightarrow+\infty.$
In particular, as stated below, strong consistency is guaranteed when $k_{n}$ is of logarithmic order.
Corollary 2.1.
Let $0<A<1/\beta$. Then, we have:
$\widehat{\beta}_{n}({A\ln n})\to\beta\text{ almost surely, as }n\rightarrow+\infty.$
Now, the following results establish the asymptotic normality of the deviation between (2) and $\beta(k_{n})$, when appropriately normalized.
Theorem 2.2 (Asymptotic normality).
Suppose that $k_{n}$ satisfies the conditions of Theorem 2.1, then
  • (i)
    Then, as $n\to+\infty$, we have the convergence in distribution:
    $\sqrt{np_{k_{n}}}\left(\widehat{\beta}_{n}({k_{n}})-\beta(k_{n})\right)\Rightarrow\mathcal{N}\left(0,\;e^{\beta}-1\right).$
  • (ii)
    In addition, asymptotic normality holds true for the ’standardized’ deviation:
    $\frac{{\sqrt{n\widehat{p}_{{k_{n}}}^{(n)}}\left({{{\widehat{\beta}}_{n}}\left({{k_{n}}}\right)-\beta\left({{k_{n}}}\right)}\right)}}{{\sqrt{{e^{{{\widehat{\beta}}_{n}}\left({{k_{n}}}\right)}}-1}}}\Rightarrow\mathcal{N}\left({0,\;1}\right),\text{ as }n\to+\infty.$
The asymptotic normality results above can be extended to the deviation between (2) and $\beta$, provided that the bias term $\beta(k_{n})-\beta$ vanishes at an appropriate rate, as stated below.
Corollary 2.2.
Suppose that the conditions of Theorem 2.2 are fulfilled. In addition, assume that $k_{n}$ is such that
$\sqrt{np_{k_{n}}}\left({1-\frac{{L\left({{e^{{k_{n}}}}}\right)}}{{L\left({{e^{{k_{n}}+1}}}\right)}}}\right)\to 0,\text{ as }n\to+\infty.$ (5)
  • (i)
    Then, we have the convergence in distribution
    $\sqrt{np_{k_{n}}}\left({\widehat{\beta}_{n}\left({{k_{n}}}\right)-\beta}\right)\Rightarrow\mathcal{N}(0,\;e^{\beta}-1)\text{ as }n\rightarrow+\infty.$
  • (ii)
    In addition, the “studentized” version is asymptotically normal:
    $\frac{{\sqrt{n\widehat{p}_{{k_{n}}}^{(n)}}\left({{{\widehat{\beta}}_{n}}(k_{n})-\beta}\right)}}{{\sqrt{{e^{{{\widehat{\beta}}_{n}}(k_{n})}}-1}}}\Rightarrow\mathcal{N}(0,1)\text{ as }n\to+\infty.$
Of course, the condition (5) on $k_{n}$ can be difficult to check in practice. This is a common issue in tail estimation and in the statistical analysis of extreme values more generally. The choice of the hyperparameter $k$ somehow governs the (asymptotic) bias-variance trade-off: the estimator (2) is expected to have a large variance when $k$ is large and to have a large bias if $k$ is too small. As depicted in Fig. 1(a), to choose $k$, one may use the same approach as that originally proposed for the Hill estimator (see e.g. [36]), which consists of plotting the values of (2) for a range of values of $k$ and choosing $k$ in a region where the estimator shows some stability. Section 3.2 of [10] suggests an adaptive algorithm for selecting $k$ for $\widehat{\beta}_{n}$, assuming the stricter condition that $F$ belongs to the class of second-order Pareto distributions.
We point out that, contrary to (variants of) the Hill inference method, the integer hyperparameter $k$ does not refer here to the number of order statistics involved in the estimator. Instead, it determines a threshold that defines the tail probability we want to estimate. This implies that the range of $k$ is the set of integers between $1$ and $\max_{1\leq i\leq n}\ln S_{i}-1$, while, for classical tail index estimators, the hyperparameter varies between $1$ and $n$.
(a) $\widehat{\beta}_{n}(k)$ estimator
(b) Hill estimator
Figure 1: (a) Behaviour of $\widehat{\beta}_{n}(k)$ for different values of $k$, to estimate the parameter $\beta=0.8$ based on a dataset of $10^{6}$ independent realizations of a Zeta distribution [^2] A discrete r.v. $W$ follows a Zeta distribution with parameter $\beta$ if $\mathbb{P}\left(W=k\right)=1/(k^{\beta-1}\zeta(\beta-1))$ where $\zeta$ is the Riemann zeta function. Its cdf satisfies $\mathbb{P}(W\geq k)\sim k^{\beta}/(\beta\zeta(\beta-1))$. This distribution is also known as Zipf’s distribution due to its relationship with Zipf’s law. . (b) Behaviour of the Hill estimator for the same dataset. Notice that the range of the x axis in both graphs is different, in the case of $\widehat{\beta}_{n}(k$ the range of $k$ is the set of integers between $1$ and $\max_{1\leq i\leq n}\ln S_{i}-1$ while in the case of the Hill method, the x-axis is the number of order statistics involved in the estimator, hence, it goes from 1 to $n$.
Remark 1.
(Quantiles vs tail probabilities) In Extreme Value Theory, it is customary to estimate the tail index by looking at the largest values, i.e. a fraction of the order statistics. The general idea consists of fixing the tail probabilities and then using the order statistics (i.e. empirical quantiles) to estimate the quantiles [12, Chapter 3]. The estimator proposed here follows the converse approach: we fix the thresholds/quantiles, next use the data to estimate the tail probabilities and finally use these estimations in order to mimic the limit behavior of (3). In general, estimators based on order statistics are difficult to analyse in a non-asymptotic way, whereas our estimate is based on probabilities of well-chosen tail events.
Remark 2.
(Averaged versions) From a practical point of view, rather than picking a single value for $k$, another natural approach would consist in averaging the estimators (2) over a range of values for the hyperparameter. Let $k$ and $m$ be such that $k>m$ and define
$\beta\left({k,m}\right)=\frac{1}{{2m+1}}\sum\limits_{j=-m}^{m}{\beta\left({k+j}\right)},\;\;\widehat{\beta}_{n}\left({k,m}\right)=\frac{1}{{2m+1}}\sum\limits_{j=-m}^{m}{\widehat{\beta}_{n}\left({k+j}\right)}.$
One may easily check that
$\beta\left({k,m}\right)=\beta+\frac{1}{{2m+1}}\left|{\ln\left({\frac{{L\left({{e^{k-m}}}\right)}}{{L\left({{e^{k+m+1}}}\right)}}}\right)}\right|.$ (6)
The non-asymptotic bound in Proposition 2.1 can be extended to the averaged version, as revealed by the analysis carried out in B in the Appendix section, as well as the strong consistency and asymptotic normality results. However, the asymptotic variance of the averaged version is shown to increase with $m$.
Remark 3.
(Second-order conditions) According to [12, Remark 3.2.6], under the following second order condition:
$\forall t>0,\alpha>0,\quad\lim_{x\to+\infty}\left({\frac{\overline{F}(tx)}{\overline{F}(x)}}-t^{\beta}\right)x^{\alpha}=0$ (7)
the Hill estimator, properly standardized, converges in distribution to a centered normal random variable. We point out that if the distribution of $S$ satisfies (7), the sequence $k_{n}$ satisfies the hypothesis of Theorem 2.2, and there is a positive constant $C$ such that $n\leq e^{Ck_{n}}$ for $n$ large enough, then the conditions of Corollary 2.2 are satisfied.
As explained in Example 1, it is possible to build a discrete heavy tail distribution that do not satisfy any second order condition (as highlighted in [28]) but for which asymptotic normality of (2) remarkably holds true.
Example 1.
Let $l\in\mathbb{N}$, $\beta>0$ and $U$ be a uniform distribution in $[0,1]$. Consider the random variable $X=10^{-l}\lfloor 10^{l}U^{-1/\beta}\rfloor$. This transformation of $U^{-1/\beta}$ turns all but the first $l$ digits behind the comma into zeros. Equation (1.6) to [28] shows that the survival function of this random variable is given by
$\bar{F}(x)=\frac{(10^{l})^{\beta}}{\left(\left\lfloor 10^{l}x\right\rfloor+1\right)^{\beta}},$ (8)
and that $\bar{F}(x)\sim x^{-\beta}$ as $x\to\infty$. Moreover, the authors showed that this distribution does not satisfy the second order condition (7). Define $L(x)={(10^{l}x)}^{\beta}({\lfloor 10^{l}x\rfloor+1})^{-\beta}$, then $L(x)\to 1$ as $x\to\infty$, hence it is slowly varying, and equation (8) can be written as $\bar{F}(x)=x^{-\beta}L(x)$. In section A.5 of the Supplementary Material we show that there exists a positive constant $K$ such that, for $x$ big enough
$\left|\frac{L(x)}{L(ex)}-1\right|\leq Kx^{-1},$ (9)
therefore, if $k_{n}\to+\infty$, for $n$ large enough, we have
$\sqrt{np_{k_{n}}}\left|{1-\frac{{L\left({{e^{{k_{n}}}}}\right)}}{{L\left({{e^{{k_{n}}+1}}}\right)}}}\right|\leq K\sqrt{\frac{nL(e^{k_{n}})}{e^{k_{n}\beta}}}\left(e^{-k_{n}}\right)\leq 2K\sqrt{ne^{-k_{n}(\beta+2)}}.$ (10)
This shows that if $k_{n}$ satisfies the conditions of Theorem 2.1 and $ne^{-k_{n}(\beta+2)}$ goes to 00, then the conditions of Corollary 2.2 are satisfied and $\sqrt{np_{k_{n}}}({\widehat{\beta}_{n}({{k_{n}}})-\beta})$ converges in distribution to a centered normal random variable. Moreover, if we take $k_{n}=A\ln n$ (as in Corollary 2.1) then, the asymptotic normality is guaranteed as long as $1/(\beta+2)<A<1/\beta$.
In the next subsection, we discuss further how the behaviour of the slowly varying function $L$ impacts the ’bias-variance’ contributions revealed by the bound (4).
2.2 Refined ’Bias vs Variance’ Analysis - Examples
We now consider several specific cases of distributions of type (1) (i.e. several instances of the slowly varying functions $L$) to explicit the asymptotic order of magnitude of the terms $1/\sqrt{np_{k+1}}$ and $|\ln(L(e^{k})/L(e^{k+1}))|$ involved in the bound (4), when $k_{n}$ is picked as in Corollary 2.1: $k_{n}=A\ln n$ with $0<A<1/\beta$.
  • The logarithmic case: Suppose that $L(n)=C\ln{n}$, where $C>0$. In this situation, we have $|\ln(L(e^{k_{n}})/L(e^{k_{n}+1}))|\sim 1/(A\ln n)$ as $n\to+\infty$, whereas $1/\sqrt{np_{k+1}}=O(1/\sqrt{n^{1-A\beta}\ln n})$.
  • The inversely logarithmic case: Consider now the situation where $L(n)=C/\ln{n}$ with $C>0$. Then, we still have we have $|\ln(L(e^{k_{n}})/L(e^{k_{n}+1}))|\sim 1/(A\ln n)$, while $1/\sqrt{np_{k+1}}=O(\sqrt{(\ln n)/n^{1-A\beta}})$ as $n\to+\infty$.
We point out that, in the two examples above, the conditions of Corollary 2.2 are not met, the bias being too big to get asymptotic normality (centered at $\beta$).
[…]
  • Slow variation with a remainder ($SR2$): Consider the case where the slowly varying function satisfies the condition $SR2$ introduced in [7]: there exist two real-valued functions $k$ and $g$ defined on $\mathbb{R}_{+}$ such that, for all $\lambda>0$,
    $\frac{L(\lambda x)}{L(x)}-1\sim\kappa(\lambda)g(x),\text{ as }x\to+\infty,$ (11)
[…]
    $L(x)=C\left(1-c|\rho|^{-1}g(x)+o\left(g\left(x\right)\right)\right),\text{ as }x\to+\infty,$ (12)
    where $C$ is a finite constant. The result below provides precise control of the bias of the estimation method in this case.
    Lemma 2.1.
    Suppose that conditions (11) and (12) are fulfilled. Then, as $n\to+\infty$, we have:
    $\ln\left(\frac{L\left(n^{A}\right)}{L\left(en^{A}\right)}\right)=-c|\rho|^{-1}n^{-A|\rho|}\left(U\left(n^{A}\right)-e^{-|\rho|}U\left(en^{A}\right)\right)+o\left(n^{-A|\rho|}U\left(n^{A}\right)\right).$
    In this situation, the bias of the method is thus of order $O(n^{-A|\rho|})$, while $1/\sqrt{np_{k+1}}$ is of order $O(n^{-(1-A\beta)/2})$. Hence, if $1/(\beta+2|\rho|)\leq A<1/\beta$, the conditions of Corollary 2.2 are satisfied with $k_{n}=A\ln n$.
To illustrate this trade-off, we present the following Monte-Carlo experiment: We generate $10^{4}$ samples of a heavy-tailed distribution and calculate $\widehat{\beta}_{10^{4}}({k})$ for all admissible values of $k$, we repeat this experiment 100 times, and then we calculate the mean and the 95% confidence interval of $\widehat{\beta}_{10^{4}}({k})$ for each value of $k$. The results of these simulations, for the cases where $L(n)$ is asymptotically constant and $L(n)$ is logarithmic, are presented in Figures 2(a) and 2(b). As expected, the behaviour of the estimator is way better in the former case than in the latter.
(a) Case when $L(n)\sim C$
(b) Case when $L(n)\sim C\ln(n)$
Figure 2: Monte-Carlo average and 95% confidence interval for $\widehat{\beta}_{10^{4}}(k)$, as $k$ is varying. The true value of $\beta$ in both cases is $0.5$.
3 Regular Null-Recurrent Chains - Regularity Index Estimation
[…]
3.1 Background and Preliminaries
Throughout the section, we suppose that the chain $X$ is $\psi$-irreducible, meaning that there exists some $\sigma$-finite measure $\psi$ on $(E,\mathcal{E})$ such that any measurable set $B\subset E$, weighted by $\psi$, can be reached by the chain with positive probability in a finite number of steps, i.e. $\sum_{n\geq 1}\Pi_{n}(x,B)>0$, no matter the starting point $x\in E$, denoting by $\Pi_{n}(x,dy)$ the $n$-th iterate of the transition probability $\Pi(x,dy)$. Recall that an irreducibility measure is said to be maximal if it dominates any other irreducibility measure. We also assume that $X$ is aperiodic (rather than replacing $\Pi$ by an iterate) and Harris recurrent, i.e. that, with probability one, it visits an infinite number of times any measurable subset $B\subset E$, weighted by maximal irreducibility measures, whatever the initial state: $\forall x\in E$, $\mathbb{P}_{x}\left(\sum_{n=1}^{\infty}\mathbb{I}\{X_{n}\in B\}=\infty\right)=1$. When Harris recurrent, a transition kernel $\Pi(x,dy)$ has a non zero invariant (positive) measure $\mu(dx)$ (i.e. such that $\int_{x\in E}\mu(dx)\Pi(x,dy)=\mu(dy)$), that is unique up to a multiplicative factor (notice incidentally that $\mu(dx)$ is a maximal irreducibility measure). Measurable sets weighted by $\mu$ are said to be Harris. For Harris recurrent chains, recall that the following strong ratio limit theorem holds. We have indeed, as $n\rightarrow\infty$,
[…]
for any initial distribution $\nu$ and any measurable sets $B$ and $C$ s.t. $\mu(C)>0$. When the measure $\mu(dx)$ is finite, the chain is said to be positive recurrent and, by convention, rather than considering $\mu(dx)/\mu(E)$, by $\mu(dx)$ we mean the stationary probability measure in this case.
Regular chains. For a wide class of Harris Markov chains, the regularity index describes how fast the occupation time related to a Harris set $B$ (i.e. the number of visits to $B$)
$\Sigma_{n}(B)=\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in B\}$
[…]
$G_{\nu,D}(t)=\frac{1}{\mu(D)}\sum_{n=1}^{\lfloor t\rfloor}\nu\Pi_{n}(D)$
where $\nu\Pi_{n}(B)=\int_{x\in E}\nu(dx)\Pi_{n}(x,B)=\mathbb{P}_{\nu}(X_{n}\in B)$ for any $B\in\mathcal{E}$. Harris recurrence entails that $G_{\nu,D}(t)\rightarrow+\infty$ as $t\rightarrow\infty$.
In the following, we restrict our attention to a specific class of Harris chains for which the rate at which $G_{\nu,D}(t)$ grows to infinity as $t\rightarrow\infty$ can be characterized. Notice that, in such cases, the rate would be independent from the pair $(\nu,D)$. Indeed, by virtue of Theorem 7.3 in [33], we have $G_{\nu_{1},D_{1}}(t)/G_{\nu_{2},D_{2}}(t)\rightarrow 1$ as $t$ goes to infinity, for any distributions $\nu_{1}$ and $\nu_{2}$ and any special sets $D_{1}$ and $D_{2}$. One may thus give the following definition, see [8, 9].
Definition 1 ($\beta$-regular Markov chain).
Let $\beta\in[0,1]$. A Harris chain $X$ is said to be $\beta$-regular if there exists a special set $D$ and a distribution $\nu$ such that the function $G_{\nu,D}$ is $\beta$-regularly varying: $\forall t>0$,
$\lim_{\lambda\rightarrow\infty}\frac{G_{\nu,D}(\lambda t)}{G_{\nu,D}(\lambda)}=t^{\beta}.$ (14)
We point out that property (14) can be rephrased as follows: there exists a slowly varying function $L_{\nu,D}(t)$ such that $G_{\nu,D}(t)=L_{\nu,D}(t)t^{\beta}$. Notice incidentally that “$\beta$-regularity” is called “$\beta$-null recurrence” in [25] when $\beta<1$, while $\beta=1$ corresponds to the positive recurrent case. The parameter $\beta$ thus rules the “frequency” at which a (supposedly regular) Harris chain $X$ recurs [8] and it is the purpose of this section to investigate the issue of estimating it with asymptotic guarantees, based on the observation of a single path $X_{1},\;\ldots,\;X_{n}$ of size $n\to+\infty$. In particular, we shall focus in subsection 3.3 on the case of regenerative chains, for which an extension of the estimator (2) can be used with statistical guarantees.
Regenerative regular chains. Recall that a Markov chain is regenerative when it possesses an accessible atom, i.e. a measurable set $A$ such that $\psi(A)>0$ and $\Pi(x,.)=\Pi(y,.)$ for all $(x,y)\in A^{2}$. By $\tau_{A}=\tau_{A}(1)=\inf\left\{n\geq 1,\;X_{n}\in A\right\}$ is meant the hitting time to $A$ and we denote by $\tau_{A}(j)=\inf\left\{n>\tau_{A}(j-1),\;X_{n}\in A\right\},$ for $j\geq 2$, the successive return times to $A$, by $\mathbb{P}_{A}$ the probability measure on the underlying space such that $X_{0}\in A$ and by $\mathbb{E}_{A}[.]$ the $\mathbb{P}_{A}$-expectation. In the regenerative case, it results from the strong Markov property that the blocks of observations in between consecutive visits to the atom
$\mathcal{B}_{1}=(X_{\tau_{A}(1)+1},...,\;X_{\tau_{A}(2)}),\;\ldots,\;\mathcal{B}_{j}=(X_{\tau_{A}(j)+1},...,\;X_{\tau_{A}(j+1)}),\;\ldots$ (15)
form a collection of i.i.d. random variables, taking their values in the torus $\mathbb{T}=\cup_{n=1}^{\infty}E^{n}$, and the sequence $\{\tau_{A}(j)\}_{j\geq 1}$, corresponding to successive times at which the chain forgets its past is a (possibly delayed) renewal process.
The class of regenerative chains includes a wide variety of Markov processes, including all chains with countable state spaces (where any recurrent state serves as an accessible atom). This class also incorporates many Markov models frequently employed in operations research and queueing theory (see, for example, [3]). Further examples of (regular) regenerative chains are presented in Examples 2 and 3 below.
Example 2.
(Bessel random walks) A Bessel random walk with drift $\delta\in[-1,\;+\infty)$ is a Markov chain with $\mathbb{N}$ as state space, jumps in $\{-1,\;+1\}$, reflecting at 00 and with transition probabilities of the form:
$\Pi(0,1)=1\quad\text{and}\quad 1-\Pi(k,k-1)=\Pi(k,k+1)=\frac{1+h(k)-\delta/(2k)}{2}\quad\forall k\geq 1,$
[…]
Example 3.
(Null recurrent, not necessarily regular, chains) By means of the model below, originally presented in [27], one can generate $\beta$-null recurrent chains for any $\beta>0$, as well as null recurrent chains that are not regular. Let $\{\eta_{n}\}_{n\in\mathbb{N}}$ be a sequence of i.i.d. real-valued random variables. Consider the chain defined by:
$X_{n}=(X_{n-1}-1)\mathbb{I}\{X_{n-1}>1\}+\eta_{n}\mathbb{I}\{X_{n-1}\in[0,1]\}\quad n\geq 1.$
This chain is regenerative, with the interval $\left[{0,1}\right]$ as atom. In addition, we have $\mathbb{P}_{x}({{\tau_{\left[{0,1}\right]}}>n})=\mathbb{P}({\lfloor{\eta_{1}}}\rfloor>n)$. Hence, $X$ is null recurrent iff $\mathbb{E}{\left\lfloor{{\eta_{1}}}\right\rfloor}=\infty$ and $\beta$-regular with $\beta\in(0,1]$ iff the r.v. $\lfloor\eta_{1}\rfloor$ has generalized discrete Pareto distribution with tail index $\beta$.
In the regenerative setting, all stochastic stability properties may be expressed in terms of speed of return to the atom. For instance, when $X$ is Harris recurrent, see Theorem 10.0.1 in [31], the invariant measure is equal to the occupation measure between two consecutive visits to the atom (up to a multiplicative factor): $\forall B\in\mathcal{E}$, $\mu(B)\sim\mathbb{E}_{A}\left[\sum_{i=1}^{\tau_{A}}\mathbb{I}\{X_{i}\in B\}\right]$. For instance, the chain is positive recurrent if and only if the expected return time to the atom is finite [^3] Its (unique) invariant probability distribution $\mu$ is then given by $\mu(B)=(1/\mathbb{E}_{A}[\tau_{A}])\mathbb{E}_{A}[\sum_{i=1}^{\tau_{A}}\mathbb{I}\{X_{i}\in B\}],\text{ for all }B\in\mathcal{E}$., i.e. $\mathbb{E}_{A}[\tau_{A}]<\infty$, see Theorem 10.2.2 in [31]. More generally, the $\beta$-regularity property can be characterized by the heaviness of the tail of the probability distribution of the regeneration times in the atomic case, as the following result shows.
Proposition 3.1.
([25], Theorem 3.1) Suppose that $X$ is regenerative Harris recurrent. Let $A$ be an atom for $X$ and $\beta\in[0,1]$. The following assertions are equivalent.
  • (i)
    The chain $X$ is $\beta$-regular.
  • (ii)
    There exists a slowly varying function $L_{A}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ such that: $\forall m\geq 1$,
    $\mathbb{P}_{A}\left(\tau_{A}\geq m\right)=L_{A}(m)\cdot m^{-\beta}.$ (16)
As a direct consequence of Proposition 3.1, we have that if $X$ is regenerative and $\beta$-regular, then $\,\beta=\sup\{\theta\in[0,1]:\;\;\mathbb{E}_{A}\left[\tau_{A}^{\theta}\right]<+\infty\}$.
Based on the decomposition (15) of the whole sample path, limit theorems for regenerative Markov chains can be derived from the application of their i.i.d. counterparts to the sequence of blocks $(\mathcal{B}_{k})_{k\geq 1}$, see e.g. [31]. This approach is usually referred to as the regenerative method and is extensively used to establish the asymptotic results stated in subsection 3.2 in the atomic case. Notice however that the regenerative blocks $\mathcal{B}_{1},\;\ldots,\;\mathcal{B}_{N_{n}}$, where $\Sigma_{n}(A)=N_{n}-1$ denotes the (random) number of regenerations before time $n$, forming the truncated trajectory up to time $n$ are not independent (the sum of their length being less than $n$ in particular), which causes technical difficulties when establishing higher-order or non-asymptotic results, see e.g. [6] or [4] and the references therein.
The inference technique for the regularity index $\beta$ of the chain $X$ developed in subsection 3.3 is based on characterization $(ii)$: the parameter $\beta$ is the tail index of a discrete generalized Pareto r.v., the regeneration time namely, i.e. the conditional survivor function of $\tau_{A}$ given $X_{0}\in A$. Incidentally, notice that the parameter $\beta$ does not depend on the atom $A$ considered (in contrast to the estimator analysed in subsection 3.3).
Based on a (random) number $N_{n}$ of (dependent) realizations of the regenerative time, namely
$S_{j}=\tau_{A}(j+1)-\tau_{A}(j)\text{ for }j=1,\;\ldots,\;N_{n},$
one may naturally compute the estimator (2). As will be shown in Theorem 3.2, in spite of the dependence structure between the $S_{j}$’s, the consistency property is preserved in the Markovian framework.
3.2 Limit Theorems of the occupation times for Regular Markov Chains
We now recall the limit results related to the behaviour of the random occupation times $\Sigma_{n}(.)$ for regular Markov chains and discuss their limitations regarding their possible use to infer the regularity index $\beta$ with (asymptotic) guarantees. The latter essentially reveals that the empirical occupation measures $\Sigma_{n}(B)$ of Harris sets $B$ grow at the sublinear rate $n^{\beta}$ (up to a slowly varying factor). As shall be seen, however, due to the great dispersion of their (asymptotic) distribution, the empirical occupation measures can hardly be used directly to estimate the key parameter $\beta$.
The result stated below claims that the logarithm of the occupation time of any Harris set provides a strongly consistent estimator of the regularity index $\beta$ when appropriately normalized. It corresponds to the comment in Remark 3.7 of [25].
Proposition 3.2.
Suppose that the chain $X$ is $\beta$-regular with $\beta\in(0,1)$ and $B$ is a Harris set. Let $\nu$ be its initial probability distribution. Then, we have:
$\widetilde{\beta}_{n}(B)\to\beta\;\;\mathbb{P}_{\nu}-\text{a.s., as }n\to+\infty,$ (17)
where
$\widetilde{\beta}_{n}(B)=\frac{\ln\Sigma_{n}(B)}{\ln n}.$ (18)
It was pointed out in Remark 3.7 of [25] that this estimator is of limited practical use due to its slow rate of convergence, although no specific rate was given therein. Equation (4.2) in [8] may suggest that $|\widetilde{\beta}_{n}(B)-\beta|$ is almost-surely $O(\ln{K_{\nu,D}(n)}/\ln n)$ as $n\to+\infty$. However, this has not been proven. In the following result, we obtain the limit distribution of the estimator and show that under the additional assumption that $L_{\nu,D}$ has a finite non-zero limit, the estimator has a logarithmic rate of convergence.
Theorem 3.1.
Suppose that the chain $X$ is $\beta$-regular with $\beta\in(0,1)$, $B$ is a Harris set, and $\nu$ is its initial probability distribution. Then, as $n\to+\infty$,
[…]
where $Z_{\beta}$ is a stable random variable with Laplace transform
$\psi_{\beta}\left(t\right)=\exp\left(-t^{\beta}/\Gamma(\beta+1)\right),\quad t\geq 0.$
In addition, if $\lim_{n\to+\infty}L_{\nu,D}(n)$ exists and is not 00, then, there exists a constant $\kappa>0$ such that
[…]
The almost sure convergence of $\widetilde{\beta}_{n}(B)$ towards $\beta$ suggests that, for $n$ large enough, $\ln\Sigma_{n}(B)\approx\beta\ln(n)$ and that the log-log plot of $\Sigma_{n}(B)$ and $n$ should look like a linear function with slope $\beta$, which could be possibly used to infer the value of $\beta$. Unfortunately, the dispersion of such a plot (and that of the process $\sigma_{n}(B)$, asymptotically described by Theorem D.1 of the Supplementary Material) is way too large in practice. To illustrate this, we simulated a Simple Symmetric Random Walk ($\beta=0.5$) with $n=10^{5}$ points, and we computed $\ln\Sigma_{\lfloor{nt}\rfloor}(B)$ for $0.1\leq t\leq 1$ (choosing $B=\{0\}$, which is an atom for this regenerative regular chain). The outcomes of this simulation are depicted in Fig. 3.
Figure 3: Log-log plot of $\ln\Sigma_{\lfloor{nt}\rfloor}(\{0\})$ in dark blue, the orange line representing the plot of the linear function $x\mapsto 0.5\times x$, while the blue area represents the 95% confidence interval for $\ln\Sigma_{\lfloor{nt}\rfloor}(\{0\})$ calculated from 100 independent trajectories.
This simulation illustrates in particular the slow convergence of $\widetilde{\beta}_{n}(B)$ described in Corollary 3.1. Hence, in the regenerative case, it is more suitable to exploit the tail behaviour of the regenerative times (cf Proposition 3.1) to estimate the regularity index $\beta$, as shall be investigated in the next subsection.
3.3 Regularity Index of a Regular Chain - Statistical Inference
Assume that $X$ is a regenerative regular chain with atom $A$ and unknown regularity index $\beta\in(0,1)$, and suppose that a sample path $X_{1},\;\ldots,\;n$ of length $n\geq 1$ is observed. Because the chain is Harris recurrent, the number of observed regeneration times $\Sigma_{n}(A)$ almost-surely tends to +$\infty$+ ∞ as $n\to+\infty$. Hence, with probability 1, we have $N_{n}\geq 1$ for $n$ large enough and one can define
$\widehat{p}^{(N_{n})}_{l}=\frac{1}{N_{n}}\sum_{i=1}^{N_{n}}\mathbb{I}\{S_{i}>e^{l}\}\text{ for any }l\in\mathbb{R}$ (19)
and form the statistic
$\widehat{\beta}_{N_{n}}(k)=\ln\left(\widehat{p}^{(N_{n})}_{k}/\widehat{p}^{(N_{n})}_{k+1}\right),$ (20)
provided that $\widehat{p}^{(N_{n})}_{k+1}>0$. We point out that, due to the randomness of $N_{n}$, (19) is a biased (strongly consistent) estimator of $p_{l}=\mathbb{P}_{A}(\tau_{A}>e^{l})$ for any $l\in\mathbb{R}$. The estimator (20) is of the same form as (2) except that the number $N_{n}$ of observations is random, and it is not independent of the sequence $\{S_{i}\}$ of observations (in particular $S_{1}+\ldots+S_{N_{n}}\leq n$).
Obtaining properties of this new estimator requires understanding the behavior of $N_{n}$, to this, the key is the following result, proved in [25], which gives the asymptotic distribution of $N_{n}$.
Proposition 3.3.
([25], Theorem 3.2) If $X$ is an atomic $\beta$-regular Markov chain with initial probability distribution $\nu$, then
[…]
where $M_{\beta}(1)$ is a Mittag-Leffler distribution with parameter $\beta$.
From Proposition 3.3 we can deduce the following corollary which provides a deterministic control of $N_{n}$ with high probability.
Corollary 3.1.
For any $\delta\in(0,1)$, let $H_{\delta}>R_{\delta}>0$ be such that
$\mathbb{P}\left(\frac{M_{\beta}(1)}{\Gamma(1-\beta)}\in\left[R_{\delta},H_{\delta}\right]\right)\geq 1-\delta/2.$
Then, there exists $n_{\delta}$ such that
$\forall n\geq n_{\delta}\quad\mathbb{P}_{\nu}\left(N_{n}\in\left[\frac{R_{\delta}}{\overline{F}(n)},\frac{H_{\delta}}{\overline{F}(n)}\right]\right)>1-\delta.$
Equipped with Corollary 3.1, we can provide the following result, which is the analogue of Proposition 2.1 in the Markovian case.
Proposition 3.4.
[…]
$\left|\widehat{\beta}_{N_{n}}(k)-\beta\right|\leq\frac{60H_{\delta}}{R_{\delta}}\sqrt{\frac{v_{n}(\delta)}{p_{k+1}}}+\left|\ln\left(\frac{L(e^{k})}{L(e^{k+1})}\right)\right|.$
Theorem 3.2.
Suppose that the atomic chain $X$ is $\beta$-regular with $\beta\in(0,1)$. Let $\nu$ be its initial probability distribution. If $k_{n}\to+\infty$ s.t. $(\ln n)\exp(k_{n}\beta)/n=o(L_{A}(\exp k_{n}))$ as $n\rightarrow+\infty$, then the estimator (20) is strongly consistent:
$\widehat{\beta}_{N_{n}}(k_{N_{n}})\to\beta\;\;\;\mathbb{P}_{\nu}-a.s\text{ as }n\to+\infty.$ (21)
In particular, strong consistency holds for $\widehat{\beta}_{N_{n}}(A\ln N_{n})$ with $A<1$.
Remark 4.
(On investigating convergence rates) Due to the impossibility of tightly controlling the sequence $N_{n}$ by a deterministic quantity in probability and the non-linearity of the estimator, we have not been able to extend to the Markovian case the asymptotic normality results of Theorem 2.2 and Corollary 2.2 via Anscombe’s theorem [18, Theorem 7.3.2]. Heuristically, if in Theorem 2.2 we take $k_{n}=\ln N_{n}$ and replace $N_{n}$ by its approximate expectation $n^{\beta}L_{\nu,A}(n)$ [25, Lemma 3.3], we would get a convergence rate of order $n^{-\beta\left(1-\beta\right)/2}L_{1}\left(n\right)$, where $L_{1}\left(n\right)$ is the slowly varying function given by $\sqrt{L_{\nu,A}(n^{\beta}L_{\nu,A}(n))/L_{\nu,A}(n)^{1-\beta}}$. This suggests a convergence rate of order $n^{-\beta(1-\beta)/2}$ when $L_{\nu,A}\sim C>0$. However, we have not been able to prove this claim.
Remark 5.
(Trajectories of random length) Suppose that the trajectory is observed until the $N$-th regeneration, i.e. $n=\tau_{A}(N)$, with $N\geq 2$. In this case, we will obtain a sequence of $N$ i.i.d blocks whose sizes follow the heavy-tailed distribution described in (1), and therefore, the results of Section 2.1 can be applied directly to this sequence. Notice that in this case, the total number of points observed in the chain (i.e. the amount of time we need to wait to collect the $N$ blocks) is a random variable that, while finite with probability one, has infinite expectation.
Remark 6.
(The (atomic) positive recurrent case) When the chain is positive recurrent (or equivalently $1$-regular), the estimator (20) can be naturally used to estimate the tail index $\beta^{\prime}\geq 1$ of the regeneration time, when the latter has a regularly varying distribution. Dedicated theoretical results can be found in section C of the Appendix.
3.4 Simulation Experiments
In order to analyze empirically the finite sample behavior of our estimator in the Markovian case, we dedicate this subsection to a simulation example. As a generative model for our experiments we use the Bessel random walks, defined in Example 2, as it will allow us to study the accuracy of our estimator, not just for different sample sizes, but also for different values of $\beta$.
For different values of $\beta$ and $n$ we have performed the following two-fold experiment:
  1. 1.
    First, we have generated one path of a $\beta$-regular Bessel random walk of length $n$. We have then computed the estimator $\widehat{\beta}_{N_{n}}(k)$ based on this trajectory and plotted the results in order to determine the value $k_{\beta,n}$ where it first stabilizes.
  2. 2.
    We have simulated $2000$ independent paths of a $\beta$-regular Bessel random walk of length $n$. Then, we have computed the estimator $\widehat{\beta}_{N_{n}}(k_{\beta,n})$ based on each of these trajectories, as well as the estimator $\widetilde{\beta}_{n}(\{0\})$.
The results of this experiment are presented in Table 1. They show that the accuracy of our estimator greatly increases as the length $n$ gets larger (notice that the bias and variance are both divided by two in order of magnitude when $n$ increases from $10^{3}$ to $10^{6}$). The histograms shown in Figure 6 hint a possible asymptotic normality of the estimator $\widehat{\beta}_{N_{n}}$. The simulation study also suggests that for moderately large sample sizes ($n\geq 10^{5}$) our estimator $\widehat{\beta}_{N_{n}}$ outperforms $\widetilde{\beta}_{n}$ as it has less bias and comparable variance, while for smaller sample sizes $\widetilde{\beta}_{n}$ works better. This behavior can be attributed to the nature of our estimator, which focuses on estimating the probabilities of tail events $\{S>e^{k}\}$ to approximate the quantity $\beta(k)$ and use the fact that, as $k$ tends to infinity, $\beta(k)$ converges to $\beta$ (see (3)). When $n$ is small, there are fewer sample points available in the tail, which forces us to select smaller values of $k$ to obtain accurate estimations of the tail probability, which worsens the estimation of $\beta$.
[…]
$\widehat{\beta}_{N_{n}}$ $\widetilde{\beta}_{n}(\{0\}$ $\widehat{\beta}_{N_{n}}$ $\widetilde{\beta}_{n}(\{0\})$
0.5 $10^{3}$ 2 0.22255 0.07987 0.11062 0.01932
$10^{4}$ 2 0.12246 0.06697 0.06032 0.01317
$10^{5}$ 3 0.01024 0.05424 0.02921 0.00825
$10^{6}$ 3 0.00343 0.04622 0.01309 0.00591
0.7 $10^{3}$ 2 0.20290 0.13154 0.12575 0.01580
$10^{4}$ 3 0.02885 0.10598 0.05673 0.00969
$10^{5}$ 3 0.00500 0.08249 0.01680 0.00604
$10^{6}$ 3 0.00476 0.06961 0.00389 0.00411
0.9 $10^{3}$ 3 0.27182 0.20158 0.15869 0.00966
$10^{4}$ 3 0.03128 0.16307 0.04685 0.00539
$10^{5}$ 4 0.02338 0.13529 0.01807 0.00286
$10^{6}$ 5 0.00704 0.11502 0.00606 0.00177
Table 1: Bias-Variance results of the simulation example.
{subcaptiongroup}
Figure 4: $\beta=0.9$, $n=10^{4}$
Figure 5: $\beta=0.5$, $n=10^{6}$
Figure 6: KDE for the estimators $\widehat{\beta}_{N_{n}}$ and $\widetilde{\beta}_{N_{n}}(\{0\})$.
3.5 Perspectives - Extension to the Pseudo-regenerative Case
Harris chains are not necessarily regenerative, of course. However, the construction proposed in [32, 33], referred to as the Nummelin splitting technique, permits to build a regenerative extension of any Harris chain (see section E for a detailed description). Essentially, given a Harris recurrent Markov chain $X$, this technique allows the construction an atomic chain, called the “split chain” that has a very similar communication structure as the original chain. Moreover, if $X$ is $\beta$-regular, then the “split chain” is also $\beta$-regular with the same value of $\beta$ [8, pp. 19].
A data-driven algorithm (see section E.1) for obtaining samples of the split chain, given a finite sample of $X$ and the transition kernel density $\pi$, has been proposed in [5]. We applied this algorithm to a simple random walk on $\mathbb{R}$, defined as $X_{0}=0$, $X_{n+1}=X_{n}+Z_{n}$ for $n\geq 1$, where $\{Z_{n}\}$ is a sequence of independent standard normal random variables. [^4] The transition kernel density of this Markov chain is $\pi(x,y)=f(y-x)$, where $f$ is the standard normal density function. After constructing the split chain, we applied our estimator to it. The results, presented in Fig. 10, show that when using the estimator $\widehat{\beta}_{N_{n}}(k)$ on the split chain, the estimates of $\beta$ closely approximate the true value ($\beta=1/2$) when $k$ is chosen near $\ln N_{n}$.
{subcaptiongroup}
Figure 7: Selection of the value of $\epsilon$.
Figure 8: $\beta$ estimations.
Figure 9: Dividing the trajectories into data blocks corresponding to the pseudo-regeneration times. The small set (-28.2203, -26.7643) is marked in red, the visits to the pseudo-atom are shown as green dotted lines.
Figure 10: Application of the pseudo-regeneration technique to construct the split chain and estimate $\beta$ using $\widehat{\beta}_{N_{n}}(k)$ in the random walk $X_{n+1}=X_{n}+Z_{n}$ where $Z_{n}$ is an i.i.d. sequence of standard normal random variables. The total number of observed points in the chain is $10^{6}$ and the number of pseudo-blocks ($N_{n}$) is 418.
When the kernel density is unknown, a procedure to approximate the split chain, based on an estimation $\widehat{\pi}_{n}$ of the kernel was presented as Algorithm 3 in [5]. As indicated in Theorem 3.1 of [6], the accuracy of this construction depends on the rate at which $\pi$ is estimated by $\widehat{\pi}_{n}$. To our knowledge, in the null-recurrent case, the sole consistent estimator of the transition density documented in the literature is the Nadaraya-Watson estimator. The proof of the consistency can be found in section 5 of [25]. However, no results regarding its rate of convergence have been established so far. Moreover, the practical choice of the bandwidth parameter involved in this estimator is a difficult and largely unresolved problem, as discussed on pp. 412 in [25]. Hence, an ambitious line of further research consists in understanding how to implement practically the approximate regenerative block construction presented in [5] in order to extend the estimation methodology studied in the previous subsection to the regular pseudo-regenerative case with theoretical guarantees.
Declarations
  • Funding: This research has been conducted as part of the project Labex MME-DII (ANR11-LBX-0023-01).
  • Competing Interests: The authors have no competing interests to declare that are relevant to the content of this article.
  • Ethics approval: Not applicable
  • Consent to participate: Not applicable
  • Consent for publication: Not applicable
  • Availability of data and materials: Not applicable
  • Code availability: The code used to generate the results in the paper is available at https://github.com/carlosds731/tail_index_estimation
  • Authors’ contributions: This paper is part of Carlos Fernández PhD thesis.
Appendix A Technical Proofs
This appendix contains the technical proofs of all the new results presented in the paper. Throughout this section we will use $u_{n}(\delta)$ to denxote $\ln\left({2/\delta}\right)/n$ for $\delta>0$ and $n\in\mathbb{N}$.
A.1 Proof of Proposition 2.1
By the triangular inequality and equation (3), we have
$\left|\widehat{\beta}_{n}(k)-\beta\right|\leq\left|\widehat{\beta}_{n}(k)-\beta(k)\right|+\left|\ln\left(\frac{L(e^{k})}{L(e^{k+1})}\right)\right|.$
Proposition 2.1 now follows by Lemma A.1.
Lemma A.1.
Let $\delta>0$ and $k$ such that $p_{k+1}\geq 16u_{n}(\delta)$, then
$\left|\widehat{\beta}_{n}(k)-\beta(k)\right|\leq 6\sqrt{\frac{u_{n}(\delta)}{p_{k+1}}},$ (22)
with probability larger than $1-2\delta$.
Proof.
In order to prove this result, we need the following lemma, proved in the Supplementary Material of [10].
Lemma A.2.
Bernstein’s inequality for Bernoulli random variables Let $W_{1},\ldots,W_{n}$ be i.i.d. samples from a distribution $F$, and we define $p_{k}=1-F(e^{k})$, ${\widehat{p}^{(n)}_{k}}=n^{-1}\sum_{i=1}^{n}{\mathbb{I}\{{{W_{i}}>{e^{k}}}\}}$. Let $\delta>0$ and take $n$ large enough so that $p_{k}\geq 4u_{n}(\delta)$. Then, with probability $1-\delta$,
$\left|\widehat{p}^{(n)}_{k}-p_{k}\right|\leq 2\sqrt{p_{k}u_{n}(\delta)}.$
Because $p_{k}\geq 16u_{n}(\delta)$ we can apply the previous lemma, then with probability greater than $1-\delta$ we have
$-2\sqrt{p_{k}u_{n}(\delta)}$ $\leq\widehat{p}^{(n)}_{k}-p_{k}\leq 2\sqrt{p_{k}u_{n}(\delta)}$
$p_{k}\left(1-2\sqrt{\frac{u_{n}(\delta)}{p_{k}}}\right)$ $\leq\widehat{p}^{(n)}_{k}\leq p_{k}\left(1+2\sqrt{\frac{u_{n}(\delta)}{p_{k}}}\right),$
taking the log in the previous equation we get
$\ln{p_{k}}+\ln{\left(1-2\sqrt{\frac{u_{n}(\delta)}{p_{k}}}\right)}$ $\leq\ln{\widehat{p}^{(n)}_{k}}\leq\ln{p_{k}}+\ln{\left(1+2\sqrt{\frac{u_{n}(\delta)}{p_{k}}}\right)}$
$\ln{\left(1-2\sqrt{\frac{u_{n}(\delta)}{p_{k}}}\right)}$ $\leq\ln{\widehat{p}^{(n)}_{k}}-\ln{p_{k}}\leq\ln{\left(1+2\sqrt{\frac{u_{n}(\delta)}{p_{k}}}\right)}$
$-3\sqrt{\frac{u_{n}(\delta)}{p_{k}}}$ $\leq\ln{\widehat{p}^{(n)}_{k}}-\ln{p_{k}}\leq 2\sqrt{\frac{u_{n}(\delta)}{p_{k}}},$ (23)
where the last pair of inequalities is obtained by using $\ln{(1+x)}\leq x$ and $\ln{(1-x)}\geq-3x/2$ for $x<1/2$. Inequality (23) implies that
$\left|\ln{\widehat{p}^{(n)}_{k}}-\ln{p_{k}}\right|\leq 3\sqrt{\frac{u_{n}(\delta)}{p_{k}}}$ (24)
with probability bigger than 1-$\delta$. Applying (24) for $k+1$ we get with probability bigger than 1-$\delta$ that
$\left|\ln{\widehat{p}^{(n)}_{k+1}}-\ln{p_{k+1}}\right|\leq 3\sqrt{\frac{u_{n}(\delta)}{p_{k+1}}}.$ (25)
Combining the triangular inequality with (24) and (25) completes the proof. ∎
A.2 Proof of Theorem 2.1
The first element in the proof of Theorem 2.1 is the following simple lemma, that shows that the non-empirical version of $\widehat{\beta}_{n}(k)$ converges to $\beta$.
Lemma A.3.
$\lim_{k\to+\infty}{\beta(k)}=\beta.$
Proof.
Because $L$ is slowly varying, $\lim_{x\to\infty}{L(tx)}/{L(x)}=1$ (see 1.2.1 of [7]) for all $t>0$, therefore ${L(e^{k})}/{L(e^{k+1})}={L(e^{k})}/{L(ee^{k})}\to 1$ and the result follows after taking limits in (3). ∎
Let $\epsilon>0$. Because $k_{n}\to\infty$, Lemma A.3 implies that $\beta(k_{n})\to\beta$, therefore we can find $N_{1}\in\mathbb{N}$ such that, for all $n\geq N_{1}$
$\left|\beta(k_{n})-\beta\right|\leq\frac{\epsilon}{2}.$ (26)
Take $\delta={2}/{n^{2}}$, then $u_{n}(\delta)={2\ln{n}}/{n}$. Because $L$ is slowly varying, $L\left(e^{k_{n}+1}\right)\sim L\left(e^{k_{n}}\right)$, then $e^{k_{n}\beta}{\ln{n}}/{n}=o\left(L\left(e^{k_{n}+1}\right)\right)$ and we can find $N_{2}\in\mathbb{N}$ such that for all for $n\geq N_{2}$ we have $p_{k_{n}+1}={L(e^{k_{n}+1})}/{e^{(k_{n}+1)\beta}}\geq{32\ln{n}}/{n}=16u_{n}(\delta)$. Therefore, we can apply Lemma A.1, obtaining that, for all $n\geq N_{2}$,
$\mathbb{P}\left(\left|\widehat{\beta}_{n}(k_{n})-\beta(k_{n})\right|\leq 6\sqrt{\frac{2\ln{n}}{np_{k_{n}+1}}}\right)\geq 1-\frac{4}{n^{2}}.$ (27)
Combining the triangular inequality with equations (26) and (27) we have that for all $n\geq\max\left(N_{1},N_{2}\right)$
$\left|\widehat{\beta}_{n}{(k_{n})}-\beta\right|\leq 6\sqrt{\frac{2\ln{n}}{np_{k_{n}+1}}}+\frac{\epsilon}{2}$ (28)
with probability bigger than $1-{4}/{n^{2}}$. Plugging $p_{k_{n}+1}={L(e^{k_{n}+1})}/{e^{(k_{n}+1)\beta}}$ in the first term of the right-hand side of (28), we get
$6\sqrt{\frac{2\ln{n}}{np_{k_{n}+1}}}=6\sqrt{\frac{2\ln{n}}{n}\frac{e^{(k_{n}+1)\beta}}{L(e^{k_{n}+1})}}=6\sqrt{2e^{\beta}}\sqrt{\frac{\ln{n}}{n}\frac{e^{(k_{n})\beta}}{L(e^{k_{n}+1})}}.$
The assumption that $e^{k_{n}\beta}{\ln{n}}/{n}=o(L(e^{k_{n}+1}))$ implies that the above equality converges to 00, therefore we can find $N_{3}\in\mathbb{N}$ such that $|6\sqrt{{2\ln{n}}/{(np_{k_{n}+1}}}|\leq{\epsilon}/{2}$ for all $n\geq N_{3}$. Then, for all $n\geq\max(N_{1},N_{2},N_{3})$
$\mathbb{P}\left(\left|\widehat{\beta}_{n}(k_{n})-\beta(k_{n})\right|\leq\epsilon\right)\geq 1-\frac{4}{n^{2}},$
which shows that $\widehat{\beta}_{n}(k_{n})$ converges in probability to $\beta$. Moreover, because the series $\sum_{n}{4}/{n^{2}}$ converges, Borell-Cantelli lemma implies that $\widehat{\beta}_{n}(k_{n})\to\beta$ almost surely.
A.3 Proof of Corollary 2.1
If we take $k_{n}=A\ln{n}$, we have $e^{\beta k_{n}}=n^{A\beta}$ then
$\lim_{n}{e^{k_{n}\beta}\frac{\ln{n}}{nL(e^{k_{n}})}}=\lim_{n}{\frac{\ln{n}}{n^{(1-A\beta)/2}}\frac{1}{n^{(1-A\beta)/2}L(n^{A})}}=0.$
For the last limit we have used that if $L$ is slowly varying, then $L(n^{A})$ is also slowly varying and that $\lim_{n}n^{\gamma}L(n)\to+\infty$ for $\gamma>0$ [7, Proposition 1.3.6.v, pp. 16]. Corollary 2.1 now follows by Theorem 2.1.
A.4 Proof of Theorem 2.2 and Corollary 2.2
Before starting with the proof of the asymptotic normality of the $\widehat{\beta}_{n}\left({k_{n}}\right)$ estimator, we need the following technical lemmas.
Lemma A.4.
Let $W_{n}$ be a sequence of positive random variables and $a_{n}$ and $b_{n}$ two positive sequences such that $a_{n}>0$, ${b_{n}}/{a_{n}}\to 0$. If there exists a random variable $W$ with continuous distribution function $F$ such that ${({W_{n}}-{a_{n}})}/{b_{n}}$ converges in distribution to $W$, then ${a_{n}}{{{(\ln{W_{n}}-\ln{a_{n}})}}/{{{b_{n}}}}}$ also converges in distribution to $W$.
Proof.
Let $x\in\mathbb{R}$ be fixed. Because ${({W_{n}}-{a_{n}})}/{b_{n}}\Rightarrow W$ in distribution, we have
$\mathbb{P}\left({{W_{n}}\leqslant{a_{n}}+{b_{n}}x}\right)\to F\left(x\right).$
Using that ${a_{n}}+{b_{n}}x={a_{n}}\left({1+a_{n}^{-1}b_{n}x}\right)$ and taking logs we get
$\mathbb{P}\left({\ln{W_{n}}\leqslant\ln{a_{n}}+\ln\left({1+\frac{{{b_{n}}}}{{{a_{n}}}}x}\right)}\right)\to F\left(x\right).$
The condition ${b_{n}}/{a_{n}}\to 0$ implies that $\ln\left({1+a_{n}^{-1}b_{n}x}\right)=a_{n}^{-1}b_{n}x+o\left(a_{n}^{-1}b_{n}\right)$. Then, $\mathbb{P}\left({{a_{n}b_{n}^{-1}}(\ln{W_{n}}-\ln{a_{n}})\leqslant x+o\left(1\right)}\right)$ converges to $F\left(x\right)$ and the Lemma follows by the continuity of $F$. ∎
Lemma A.5.
If $k_{n}$ satisfies the hypothesis of Theorem 2.1, then,
$\frac{{\widehat{p}_{{k_{n}}}^{n}}}{{\bar{F}\left({{e^{{k_{n}}}}}\right)}}\to 1\quad\textit{almost surely}.$
Proof.
By Lemma A.2, for any $\delta>0$ such that $p_{k}\geq 4u_{n}\left(\delta\right)$ we have that,
$\mathbb{P}\left({\left|{\frac{{\widehat{p}_{k}^{n}}}{{{p_{k}}}}-1}\right|\leqslant 2\sqrt{\frac{{{u_{n}}\left(\delta\right)}}{{{p_{k}}}}}}\right)\geqslant 1-\delta.$ (29)
As in the proof of Theorem 2.1, let $\delta={2}/{n^{2}}$, so $u_{n}(\delta)={2\ln{n}}/{n}$. The condition $e^{k_{n}\beta}{\ln{n}}/{n}=o\left({L\left({{e^{{k_{n}}}}}\right)}\right)$ implies that we can find $N_{1}\in\mathbb{N}$ such that ${p_{{k_{n}}}}\geqslant{{8\ln n}}/{n}$ for all $n\geq N_{1}$, therefore, by equation (29), we have, for all $n\geqslant{N_{1}}$
$\mathbb{P}\left({\left|{\frac{{\widehat{p}_{{k_{n}}}^{n}}}{{{p_{{k_{n}}}}}}-1}\right|\leqslant 2\sqrt{\frac{{2\ln n}}{{n{p_{{k_{n}}}}}}}}\right)\geqslant 1-\frac{2}{{{n^{2}}}}.$
Let $\epsilon>0$. Notice that ${{\ln n}}{{(n{p_{{k_{n}}}})^{-1}}}={{{e^{{k_{n}}\beta}}\ln n}}/{(nL\left({{e^{{k_{n}}}}}\right))}$ and this goes to 0 as $n$ goes to +$\infty$+ ∞, therefore, we can find $N_{2}$ such that $2\sqrt{{{2\ln n}}/{{(n{p_{{k_{n}}}})}}}\leqslant\epsilon$ for all $n\geq N_{2}$, then, for all $n\geqslant\max\left({{N_{1}},{N_{2}}}\right)$
$\mathbb{P}\left({\left|{\frac{{\widehat{p}_{{k_{n}}}^{n}}}{{{p_{{k_{n}}}}}}-1}\right|\leqslant\epsilon}\right)\geqslant 1-\frac{2}{{{n^{2}}}},$
and the result follows by Borel-Cantelli’s Lemma. ∎
The next result can be obtained using the same arguments of Example 3.11 on [13].
Lemma A.6.
Let $W_{n}$ be a sequence of i.i.d. random variables with survival function (1), $\phi_{1}$ and $\phi_{2}$ bounded functions and $u_{n}$ an increasing sequence of real numbers such that $u_{n}\to+\infty$. Define
${W_{n,i}}=\frac{{{W_{i}}}}{{{u_{n}}}}\mathbb{I}\left\{{\frac{{{W_{i}}}}{{{u_{n}}}}>1}\right\},\quad{v_{n}}=\mathbb{P}\left({{W_{n,i}}\neq 0}\right)\quad\textnormal{and}$
${\widetilde{Z}_{n}}\left({{\phi_{k}}}\right)=\frac{1}{{\sqrt{n{v_{n}}}}}\sum\limits_{i=1}^{n}{\left({{\phi_{k}}\left({{W_{n,i}}}\right)-\mathbb{E}{\phi_{k}}\left({{W_{n,i}}}\right)}\right)}.$
If the following conditions hold:
  • (A1)
    $nv_{n}\to+\infty$,
  • (A2)
    $\mathbb{E}{\left[{{{\phi_{k}}\left({{W_{n,1}}}\right)}^{4}}\right]}=O\left({{v_{n}}}\right),\quad,k=1,2,$
  • (A3)
    $\mathop{\lim}\limits_{n}\frac{1}{{{v_{n}}}}{{\mathbb{E}\left[{{\phi_{k}}\left({{W_{n,1}}}\right){\phi_{l}}\left({{W_{n,1}}}\right)}\right]}}={\sigma_{kl}},\quad k,l\in\{1,2\},$
then ${\left({{{\widetilde{Z}}_{n}}\left({{\phi_{k}}}\right)}\right)_{1\leqslant k\leqslant 2}}$ converges weakly to a centred normal distribution with covariance matrix ${\left({{\sigma_{kl}}}\right)_{1\leqslant k,l\leqslant 2}}$.
Let $k_{n}$ satisfy the conditions of Theorem 2.1, take $u_{n}=e^{k_{n}}$, ${\phi_{1}}\left(x\right)=\mathbb{I}\left\{{x>1}\right\}$ and ${\phi_{2}}\left(x\right)=\mathbb{I}\left\{{x>e}\right\}$. With this notation:
${\phi_{1}}\left({{W_{n,i}}}\right)$ $=\mathbb{I}\left\{{\left({\frac{{{W_{i}}}}{{{u_{n}}}}\mathbb{I}\left\{{\frac{{{W_{i}}}}{{{u_{n}}}}>1}\right\}}\right)>1}\right\}=\mathbb{I}\left\{{\frac{{{W_{i}}}}{{{u_{n}}}}>1}\right\},$
${\phi_{2}}\left({{W_{n,i}}}\right)$ $=\mathbb{I}\left\{{\left({\frac{{{W_{i}}}}{{{u_{n}}}}\mathbb{I}\left\{{\frac{{{W_{i}}}}{{{u_{n}}}}>1}\right\}}\right)>e}\right\}=\mathbb{I}\left\{{\frac{{{W_{i}}}}{{{u_{n}}}}>e}\right\},$
$\mathbb{E}\left[{\phi_{1}}\left({{W_{n,i}}}\right)\right]$ $=\mathbb{P}\left({\frac{{{W_{i}}}}{{{u_{n}}}}>1}\right)=\overline{F}\left({{u_{n}}}\right),$
$\mathbb{E}\left[{\phi_{2}}\left({{W_{n,i}}}\right)\right]$ $=\mathbb{P}\left({\frac{{{W_{i}}}}{{{u_{n}}}}>e}\right)=\overline{F}\left({e{u_{n}}}\right),$
${v_{n}}$ $=\mathbb{P}\left({{W_{n,i}}\neq 0}\right)=\mathbb{P}\left({\frac{{{W_{i}}}}{{{u_{n}}}}>1}\right)=\overline{F}\left({{u_{n}}}\right).$
Let $w_{n}=\overline{F}\left({e{u_{n}}}\right)$, ${\lambda_{n}}={{\overline{F}\left({{u_{n}}}\right)}}/{{\overline{F}\left({e{u_{n}}}\right)}}={v_{n}}/{w_{n}}$ (notice that $\lambda_{n}\to e^{\beta}$) and ${y_{n}}=\sqrt{{{{v_{n}}}}/{{(n{w_{n}}^{2})}}}$, then,
$\widehat{\lambda}_{n}$ $=\frac{{\sum\limits_{i=1}^{n}{\mathbb{I}\left\{{{W_{i}}>u_{n}}\right\}}}}{{\sum\limits_{i=1}^{n}{\mathbb{I}\left\{{{W_{i}}>eu_{n}}\right\}}}}=\frac{{n\mathbb{E}\left[{\phi_{1}}\left({{W_{n,1}}}\right)\right]+\sum\limits_{i=1}^{n}{\left\{{{\phi_{1}}\left({{W_{n,i}}}\right)-\mathbb{E}\left[{\phi_{1}}\left({{W_{n,1}}}\right)\right]}\right\}}}}{{n\mathbb{E}\left[{\phi_{2}}\left({{W_{n,1}}}\right)\right]+\sum\limits_{i=1}^{n}{\left\{{{\phi_{2}}\left({{W_{n,i}}}\right)-\mathbb{E}\left[{\phi_{2}}\left({{W_{n,1}}}\right)\right]}\right\}}}}$
$=\frac{{\frac{{\mathbb{E}\left[{\phi_{1}}\left({{W_{n,1}}}\right)\right]}}{{\mathbb{E}\left[{\phi_{2}}\left({{W_{n,1}}}\right)\right]}}+\frac{{\sum\limits_{i=1}^{n}{\left\{{{\phi_{1}}\left({{W_{n,i}}}\right)-\mathbb{E}\left[{\phi_{1}}\left({{W_{n,1}}}\right)\right]}\right\}}}}{{n\mathbb{E}\left[{\phi_{2}}\left({{W_{n,1}}}\right)\right]}}}}{{1+\frac{{\sum\limits_{i=1}^{n}{\left\{{{\phi_{2}}\left({{W_{n,i}}}\right)-\mathbb{E}\left[{\phi_{2}}\left({{W_{n,1}}}\right)\right]}\right\}}}}{{n\mathbb{E}\left[{\phi_{2}}\left({{W_{n,1}}}\right)\right]}}}}=\frac{{\frac{{\mathbb{E}\left[{\phi_{1}}\left({{W_{n,1}}}\right)\right]}}{{\mathbb{E}\left[{\phi_{2}}\left({{W_{n,1}}}\right)\right]}}+{{\widetilde{Z}}_{n}}\left({{\phi_{1}}}\right)\sqrt{\frac{{{v_{n}}}}{{n{w_{n}}^{2}}}}}}{{1+{{\widetilde{Z}}_{n}}\left({{\phi_{2}}}\right)\sqrt{\frac{{{v_{n}}}}{{n{w_{n}}^{2}}}}}}$
$=\frac{{{\lambda_{n}}+{{\widetilde{Z}}_{n}}\left({{\phi_{1}}}\right){y_{n}}}}{{1+{{\widetilde{Z}}_{n}}\left({{\phi_{2}}}\right){y_{n}}}}.$ (30)
In order to apply Lemma A.6, we need to check its hypotheses. We will start by (A1). By hypothesis, $e^{k_{n}\beta}{\ln{n}}/{n}=o\left(L\left(e^{k_{n}+1}\right)\right)$ and $L$ is slowly varying, hence, we can write ${e^{{k_{n}}\beta}}{{\ln n}}/{n}=L\left({{e^{{k_{n}}}}}\right)\varepsilon\left(n\right)$ where $\varepsilon\left(n\right)\to 0$, then
$nv_{n}=n\overline{F}({{u_{n}}})=n\frac{L(e^{k_{n}})}{e^{k_{n}\beta}}=\frac{\ln n}{\epsilon(n)}\to+\infty.$
Hypotheses (A2) follows directly from the following calculations
$\mathbb{E}\left[\phi_{1}(W_{n,1})^{4}\right]$ $=\mathbb{E}\left[\phi_{1}(W_{n,1})\right]=\mathbb{P}\left(W_{i}>u_{n}\right)=\overline{F}(u_{n})=v_{n},$
$\mathbb{E}\left[\phi_{2}(W_{n,1})^{4}\right]$ $=\mathbb{E}\left[\phi_{2}(W_{n,1})\right]=\mathbb{P}\left(W_{i}>eu_{n}\right)=\overline{F}(eu_{n})\leq v_{n}.$
Finally, for (A3), observe that
$\mathbb{E}\left[{{\phi_{1}}\left({{W_{n,1}}}\right)^{2}}\right]$ $=v_{n},$
$\mathbb{E}\left[{{\phi_{1}}\left({{W_{n,1}}}\right){\phi_{2}}\left({{W_{n,1}}}\right)}\right]$ $=w_{n},$
$\mathbb{E}\left[{{\phi_{2}}\left({{W_{n,1}}}\right)^{2}}\right]$ $=w_{n},$
therefore, the limits stated in (A3) exist and their values are $\sigma_{11}=1$, $\sigma_{12}=\sigma_{22}=e^{-\beta}$.
Applying Lemma A.6 we obtain that ${({\widetilde{{Z}}_{n}\left({{\phi_{k}}}\right)})_{1\leqslant k\leqslant 2}}$ converges to a centred normal distribution with covariance matrix ${\left({{\sigma_{kl}}}\right)_{1\leqslant k,l\leqslant 2}}$. Taking into account that ${y_{n}}\sim{{{e^{\beta}}}}/{{\sqrt{n{v_{n}}}}}$, it follows that
${\widehat{\lambda}_{n}}$ $=\left({{\lambda_{n}}+{{\widetilde{Z}}_{n}}\left({{\phi_{1}}}\right){y_{n}}}\right)\left({1-{{\widetilde{Z}}_{n}}\left({{\phi_{2}}}\right){y_{n}}+o_{P}\left({\frac{1}{{\sqrt{n{v_{n}}}}}}\right)}\right)$
$={\lambda_{n}}+{y_{n}}\left({{{\widetilde{Z}}_{n}}\left({{\phi_{1}}}\right)-{\lambda_{n}}{{\widetilde{Z}}_{n}}\left({{\phi_{2}}}\right)}\right)+{o_{P}}\left({\frac{1}{{\sqrt{n{v_{n}}}}}}\right).$ (31)
Then, $\sqrt{n{v_{n}}}({{{\widehat{\lambda}}_{n}}-{\lambda_{n}}})$ converges weakly to a centred normal distribution with variance ${e^{2\beta}}\left({{\sigma_{11}}+{e^{2\beta}}{\sigma_{22}}-2{e^{\beta}}{\sigma_{12}}}\right)={e^{2\beta}}\left({{e^{\beta}}-1}\right)$. This can be resumed in the following lemma.
Lemma A.7.
Let $W_{n}$ and $u_{n}$ be as in Lemma A.6, if $k_{n}$ satisfies the conditions of Theorem 2.1, then the following convergence in distribution holds
$\sqrt{n\bar{F}\left({{e^{{k_{n}}}}}\right)}\left({\frac{{\sum\limits_{i=1}^{n}{\mathbb{I}\left\{{{W_{i}}>{e^{{k_{n}}}}}\right\}}}}{{\sum\limits_{i=1}^{n}{\mathbb{I}\left\{{{W_{i}}>{e^{{k_{n}}+1}}}\right\}}}}-\frac{{\bar{F}\left({{e^{{k_{n}}}}}\right)}}{{\bar{F}\left({{e^{{k_{n}}+1}}}\right)}}}\right)\Rightarrow\mathcal{N}\left(0,\;e^{2\beta}(e^{\beta}-1)\right).$
Lemmas A.3, A.4 and A.7 combined with equation (3) imply the first part of Theorem 2.2, the second part follows from Lemma A.5 and Slutsky’s Theorem. Corollary 2.2 follows immediately.
A.5 Technical details of Example 1
To establish equation (9), we proceed as follows. For any $x>1$, we can express:
$\frac{L(x)}{L(ex)}=\left(\frac{10^{l}x}{\lfloor 10^{l}x\rfloor+1}\frac{\lfloor 10^{l}ex\rfloor+1}{10^{l}ex}\right)^{\beta}=e^{-\beta}\left(\frac{\lfloor 10^{l}ex\rfloor+1}{\lfloor 10^{l}x\rfloor+1}\right)^{\beta}.$
Using that $y-1<\lfloor y\rfloor\leq y$ for any $y>0$, we can derive:
$\frac{\lfloor 10^{l}ex\rfloor+1}{\lfloor 10^{l}x\rfloor+1}\leq\frac{10^{l}ex+1}{10^{l}x}\leq e+\frac{1}{10^{l}x},$
consequently,
$\frac{L(x)}{L(ex)}\leq\left(1+\frac{1}{10^{l}ex}\right)^{\beta}=1+\frac{\beta}{10^{l}ex}+o(x^{-1}).$
By a similar argument, if follows that:
$\frac{L(x)}{L(ex)}\geq\left(1-\frac{1}{10^{l}x+1}\right)^{\beta}=1-\frac{\beta}{10^{l}x+1}+o(x^{-1}).$
As a result, for sufficiently large $x$, we obtain that
$\left|\frac{L(x)}{L(ex)}-1\right|\leq\frac{2\beta}{10^{l}}x^{-1}.$
A.6 Proof of Lemma 2.1
The representation is a direct application of Lemma A.8 and the fact that $g$ is a regularly varying function of index $\rho$.
Lemma A.8.
Assume that $L$ satisfies $SR2$, has positive decrease and $x$ is big enough such that representation the (12) holds, then
$\ln\left(\frac{L(x)}{L(\lambda x)}\right)=-c|\rho|^{-1}\left(g(x)-g(\lambda x)\right)+o\left(g(x)\right).$ (32)
Proof.
Denote $A(x)=c\rho^{-1}g(x)+o\left(g\left(x\right)\right)$. By (12) we have
$\ln\left(\frac{L(x)}{L(\lambda x)}\right)$ $=\ln\left(\frac{C(1+A(x))}{C(1+A(\lambda x))}\right)=\ln\left(\frac{1+A(x)}{1+A(\lambda x)}\right)$
$=\ln{\left(1+A\left(x\right)\right)}-\ln(1+A(\lambda x)).$ (33)
Using the first order expansion for $\ln(1+A(x))$ we have that
$\ln\left(1+A\left(x\right)\right)$ $=c\rho^{-1}g(x)+o\left(g\left(x\right)\right)+\underbrace{o\left(c\rho^{-1}g(x)+o\left(g\left(x\right)\right)\right)}_{o\left(g\left(x\right)\right)}$
$=c\rho^{-1}g(x)+o\left(g\left(x\right)\right).$ (34)
Applying (34) to $\lambda x$ we get
$\ln\left(1+A\left(\lambda x\right)\right)=c\rho^{-1}g(\lambda x)+o\left(g\left(x\right)\right),$ (35)
where we have used that if $g$ is regularly varying then $o\left(g\left(\lambda x\right)\right)=o\left(g\left(x\right)\right)$. The result now follows by plugging (34) and (35) into (33). ∎
A.7 Proof of Theorem 3.1
Theorem 3.1 is the the result of the Continuous Mapping Theorem and the following Theorem, which is the particularization of Theorem 2.3 in [8] to the indicator function $\mathbb{I}_{B}$. In section D, we provide a functional generalization of Theorem A.1.
Theorem A.1 (Limit distributions).
Let $\beta\in[0,1)$ and $\nu$ be any probability distribution on $E$. Suppose that the chain $X$ is $\beta$-regular and let $B$ be a Harris set with finite and strictly positive $\mu$-measure. We have the following convergences in $\mathbb{P}_{\nu}$-distribution.
  • (i)
    If $\beta=0$, we then have, as $n\rightarrow\infty$,
[…]
    where $\mathcal{E}(\lambda)$ denotes the exponential distribution with mean $01/\lambda>$ 0.
  • (ii)
    If $\beta\in(0,1)$, we have, as $n\rightarrow\infty$,
[…]
    where $Z_{\beta}$ is a stable random variable with Laplace transform
    $\psi_{\beta}\left(t\right)=\exp\left(-t^{\beta}/\Gamma(\beta+1)\right),\quad t\geq 0.$
A.8 Proof of Corollary 3.1
The random variable $\frac{M_{\beta}(1)}{\Gamma(1-\beta)}$ is continuous and positive, therefore, we can find positive constants $R_{\delta}$ and $H_{\delta}$ such that
$\mathbb{P}\left({\frac{M_{\beta}(1)}{\Gamma(1-\beta)}\in\left[{R_{\eta},H_{\eta}}\right]}\right)\geqslant 1-\frac{\delta}{2}.$ (36)
By Proposition 3.3, $N_{n}\overline{F}(n)$ converges in distribution to $\frac{M_{\beta}(1)}{\Gamma(1-\beta)}$, hence, there exists $n_{\delta}$ such that
$\left|\mathbb{P}\left(N_{n}\overline{F}(n)\in\left[{R_{\eta},H_{\eta}}\right]\right)-\mathbb{P}\left({\frac{M_{\beta}(1)}{\Gamma(1-\beta)}\in\left[{R_{\eta},H_{\eta}}\right]}\right)\right|<\frac{\delta}{2}\quad\forall n\geq n_{\delta}.$ (37)
The result now follows by combining (36) and (37).
A.9 Proof of Proposition 3.4
As in Proposition 2.1, the result is a direct consequence of the Lemma A.10 and the inequality
$\left|\widehat{\beta}_{N_{n}}(k)-\beta\right|\leq\left|\widehat{\beta}_{N_{n}}(k)-\beta(k)\right|+\left|\ln\left(\frac{L(e^{k})}{L(e^{k+1})}\right)\right|.$
Lemma A.9.
If $0<t<p_{k}$ then,
$\mathbb{P}\left(\left|\widehat{p}^{(n)}_{k}-p_{k}\right|>t\right)\leq 2\exp{\left(-\frac{nt^{2}}{4p_{k}}\right)}$ (38)
Proof.
By Bernstein’s inequality [38, Lemma 19.32], for all $y>0$
$\mathbb{P}\left(\sqrt{n}\left|\widehat{p}^{(n)}_{k}-p_{k}\right|>y\right)$ $\leq 2\exp{\left(-\frac{1}{4}\frac{y^{2}}{p_{k}+t/\sqrt{n}}\right)}$
$\leq 2\min\left(\exp\left(-\frac{1}{4}\frac{y^{2}}{p_{k}}\right),\exp\left(-\frac{1}{4}\sqrt{n}{p_{k}}\right)\right),$
and if $y<\sqrt{n}p_{k}$, then
$\mathbb{P}\left(\sqrt{n}\left|\widehat{p}^{(n)}_{k}-p_{k}\right|>y\right)\leq 2\exp\left(-\frac{1}{4}\frac{y^{2}}{p_{k}}\right).$
The result now follows by setting $y=t\sqrt{n}$. ∎
Lemma A.10.
Let $\delta>0$ and $n\geq n_{\delta}$ such that $p_{k+1}\geq{({40H_{\delta}}R^{-1}_{\delta})}^{2}v_{n}(\delta)$, then
$\mathbb{P}_{\nu}\left(\left|\widehat{\beta}_{N_{n}}(k)-\beta(k)\right|\leq\frac{60H_{\delta}}{R_{\delta}}\sqrt{\frac{v_{n}(\delta)}{p_{k+1}}}\right)>1-2\delta,$ (39)
where $H_{\delta}$, $R_{\delta}$ and $n_{\delta}$ are as in Corollary 3.1.
Proof.
Let $\delta>0$ be fixed. With $R_{\delta}$ and $H_{\delta}$ as in Corollary 3.1, define the events
$\mathcal{E}_{n}=\left\{N_{n}\in\left[\frac{R_{\delta}}{\overline{F}(n)},\frac{H_{\delta}}{\overline{F}(n)}\right]\right\},$
and denote by ¯$\overline{\mathcal{E}}_{n}$ the complement of $\mathcal{E}_{n}$. By Corollary 3.1, there exists $n_{\delta}$ such that (¯$\mathbb{P}(\overline{\mathcal{E}}_{n})<\delta/2$. Let $t>0$, then, for each $n\geq n_{\delta}$ it holds that
$\mathbb{P}\left(\left|\widehat{p}^{(N_{n})}_{k}-p_{k}\right|>t\right)$ ($\leqslant\mathbb{P}\left(\left\{\left|\widehat{p}^{(N_{n})}_{k}-p_{k}\right|>t\right\}\cap\mathcal{E}_{n}\right)+\mathbb{P}\left(\overline{\mathcal{E}}_{n}\right),$
($\leqslant\mathbb{P}\left(\left\{\left|\widehat{p}^{(N_{n})}_{k}-p_{k}\right|>t\right\}\cap\mathcal{E}_{n}\right)+\frac{\delta}{2}.$
Notice that, on $\mathcal{E}_{n}$ we have the following inclusions
$\left\{\left|\widehat{p}^{(N_{n})}_{k}-p_{k}\right|>t\right\}$ =$=\left\{\frac{1}{N_{n}}\left|\sum_{i=1}^{N_{n}}{\mathbb{I}\left\{S_{i}>e^{k}\right\}}-p_{k}\right|>t\right\}$
=$=\left\{\left|\sum_{i=1}^{N_{n}}{\mathbb{I}\left\{S_{i}>e^{k}\right\}}-p_{k}\right|>N_{n}t\right\}$
$\subseteq\left\{\max_{1\leq m\leq\frac{H_{\delta}}{\overline{F}(n)}}\left|\sum_{i=1}^{m}{\mathbb{I}\left\{S_{i}>e^{k}\right\}}-p_{k}\right|>\frac{R_{\delta}t}{\overline{F}(n)}\right\},$
therefore,
($\mathbb{P}\left(\left\{\left|\widehat{p}^{(N_{n})}_{k}-p_{k}\right|>t\right\}\cap\mathcal{E}_{n}\right)\leqslant\mathbb{P}\left(\max_{1\leq m\leq\frac{H_{\delta}}{\overline{F}(n)}}\left|\sum_{i=1}^{m}{\mathbb{I}\left\{S_{i}>e^{k}\right\}}-p_{k}\right|>\frac{R_{\delta}t}{\overline{F}(n)}\right).$
By Montgomery-Smith’s inequality [29, Corollary 4], the right-hand side of the previous inequality is smaller than
$9\mathbb{P}\left(\left|\sum_{i=1}^{\frac{H_{\delta}}{\overline{F}(n)}}{\mathbb{I}\left\{S_{i}>e^{k}\right\}}-p_{k}\right|>\frac{R_{\delta}t}{10\overline{F}(n)}\right)=9\mathbb{P}\left(\left|\widehat{p}^{\left(\frac{H_{\delta}}{\overline{F}(n)}\right)}_{k}-p_{k}\right|>\frac{R_{\delta}t}{10H_{\delta}}\right)$
Using Lemma A.9 we obtain that, for $t<\frac{10p_{k}H_{\delta}}{R_{\delta}}$ it holds that
$\mathbb{P}\left(\left|\widehat{p}^{\left(\frac{H_{\delta}}{\overline{F}(n)}\right)}_{k}-p_{k}\right|>\frac{R_{\delta}t}{10H_{\delta}}\right)\leq 2\exp\left(-\frac{R^{2}_{\delta}}{400H_{\delta}}\frac{t^{2}}{\overline{F}(n)p_{k}}\right).$
Therefore for all $t<\frac{10p_{k}H_{\delta}}{R_{\delta}}$ we have that
$\mathbb{P}\left(\left|\widehat{p}^{(N_{n})}_{k}-p_{k}\right|>t\right)\leq 18\exp\left(-\frac{R^{2}_{\delta}}{400H_{\delta}}\frac{t^{2}}{\overline{F}(n)p_{k}}\right)+\frac{\delta}{2}$ (40)
Taking $t=20R^{-1}_{\delta}\sqrt{\ln{(36/\delta)}\overline{F}(n)p_{k}H_{\delta}}=20H_{\delta}R^{-1}_{\delta}\sqrt{v_{n}(\delta)p_{k}}$ in (41) implies that, if $n$ is large enough such that $p_{k}>4v_{n}(\delta)$ and $n\geq n_{\delta}$, then
$\mathbb{P}\left(\left|\widehat{p}^{(N_{n})}_{k}-p_{k}\right|>20\frac{H_{\delta}}{R_{\delta}}\sqrt{v_{n}p_{k}}\right)\leq\delta.$ (41)
Equation (39) now follows by the same argument used in the proof of Proposition 2.1. ∎
A.10 Proof of Theorem 3.2
The recurrence of the chain implies that $N_{n}$ converges almost surely to +$\infty$+ ∞, then, by Theorem 8.1 in page 302 of [18], we can replace $n$ by $N_{n}$ on the strong consistency results we presented on section 2, to obtain equivalent results for the sequence $S_{1}\ldots,S_{N_{n}}$.
Appendix B Averaged Estimators
Here we collect some remarks and results related to the averaged estimator $\widehat{\beta}_{n}\left({k,m}\right)$. First, we detail how to get the expression (6) from (3). Let $k>0$ be fixed, for each $j$ we have:
$\beta\left({k+j}\right)=\beta+\ln\left({\frac{{L\left({{e^{k+j}}}\right)}}{{L\left({{e^{k+j+1}}}\right)}}}\right),$
then,
$\frac{1}{{2m+1}}\sum\limits_{j=-m}^{m}{\beta\left({k+j}\right)}$ $=\frac{1}{{2m+1}}\sum\limits_{j=-m}^{m}\beta+\frac{1}{{2m+1}}\sum\limits_{j=-m}^{m}{\ln\left({\frac{{L\left({{e^{k+j}}}\right)}}{{L\left({{e^{k+j+1}}}\right)}}}\right)}$
$=\beta+\frac{1}{{2m+1}}\ln\left({\prod\limits_{j=-m}^{m}{\frac{{L\left({{e^{k+j}}}\right)}}{{L\left({{e^{k+j+1}}}\right)}}}}\right)$
$=\beta+\frac{1}{{2m+1}}\ln\left({\frac{{L\left({{e^{k-m}}}\right)}}{{L\left({{e^{k+m+1}}}\right)}}}\right).$
Our first result in this regard is the concentration inequality equivalent to (4) but for the averaged version of the estimator.
Proposition B.1.
[…]
$\left|{\widehat{\beta}_{n}\left({k,m}\right)-\beta}\right|\leqslant 6\sqrt{\frac{{{u_{n}}\left(\delta\right)}}{{{p_{k+m+1}}}}}+\frac{1}{{2m+1}}\left|{\ln\left({\frac{{L\left({{e^{k-m}}}\right)}}{{L\left({{e^{k+m+1}}}\right)}}}\right)}\right|.$ (42)
The following results show that, for well-chosen $k_{n}$ and $m_{n}$, the estimator $\widehat{\beta}_{n}(k_{n},m_{n})$ is strongly consistent.
Theorem B.1 (Strong consistency).
Let $k_{n}$ and $m_{n}$ such that, as $n\to\infty$
  • i)
    ${k_{n}}-{m_{n}}\to+\infty$,
  • ii)
    $\sum\limits_{n=1}^{+\infty}{\frac{4}{{{n^{2}}}}\left({1-2{m_{n}}}\right)}$ is convergent,
  • iii)
    ${e^{\left({{k_{n}}+{m_{n}}}\right)\beta}}\frac{{\ln n}}{n}=o\left({L\left({{e^{{k_{n}}+{m_{n}}}}}\right)}\right);$
then, $\widehat{\beta}_{n}\left({k_{n},m_{n}}\right)$ converges almost surely to $\beta$.
Corollary B.1.
Let $A,l$ be a positive numbers such that $l>1$ and $l>{A\beta}/{(1-A\beta)}$ then
$\widehat{\beta}_{n}\left({A\ln n,\frac{{A\ln n}}{l}}\right)\to\beta\quad a.s.$
B.1 Proof of Proposition B.1
Similarly to the proof of Theorem 2.1, Theorem B.1 follows by triangular inequality, the definition of $\beta\left({k,m}\right)$ and the following Lemma B.1, which provides us a bound for the difference between $\widehat{\beta}(k,m)-\beta(k,m)$.
Lemma B.1.
Let $\delta>0$ and $k$ and $m$ such that $p_{k+m+1}\geq 16u_{n}(\delta)$, then
$\left|{\widehat{\beta}_{n}(k,m)-\beta(k,m)}\right|\leqslant 6\sqrt{\frac{{{u_{n}}(\delta)}}{{{p_{k+m+1}}}}},$ (43)
with probability larger than $1-2\delta\left({1-2m}\right)$.
Proof.
For the left-hand side of equation (43) we have
$\left|{\widehat{\beta}_{n}(k,m)-\beta(k,m)}\right|$ $=\frac{1}{{2m+1}}\left|{\sum\limits_{j=-m}^{m}{\left({\widehat{\beta}_{n}\left({k+j}\right)-\beta\left({k+j}\right)}\right)}}\right|$
$=\frac{1}{{2m+1}}\left|{\sum\limits_{j=0}^{2m}{\left({\widehat{\beta}_{n}\left({k-m+j}\right)-\beta\left({k-m+j}\right)}\right)}}\right|,$
then,
$\left|{\widehat{\beta}_{n}(k,m)-\beta(k,m)}\right|\leqslant\frac{1}{{2m+1}}\sum\limits_{j=0}^{2m}{\left|{\widehat{\beta}_{n}\left({k-m+j}\right)-\beta\left({k-m+j}\right)}\right|}.$ (44)
Given $p_{k+m+1}\geq 16u_{n}(\delta)$, it follows that ${p_{k-m+j+1}}\geqslant 16{u_{n}}(\delta)$ for all $j$ between 0 and $2m$, allowing the application of Lemma A.1. This yields, for each $j$,
$\left|{\widehat{\beta}_{n}\left({k-m+j}\right)-\beta\left({k-m+j}\right)}\right|\leqslant 6\sqrt{\frac{{{u_{n}}(\delta)}}{{{p_{k-m+j+1}}}}},$ (45)
with probability bigger than $1-2\delta$.
The joint probability of inequality (45) holding for all $j$ between 00 and $2m$ is greater than $\left({2m+1}\right)\left({1-2\delta}\right)-2m=1-2\delta\left({1-2m}\right)$. Hence, with at least this probability, equation (44) simplifies to,
$\left|{\widehat{\beta}_{n}(k,m)-\beta(k,m)}\right|\leqslant\frac{1}{{2m+1}}\sum\limits_{j=0}^{2m}{6\sqrt{\frac{{{u_{n}}(\delta)}}{{{p_{k-m+j+1}}}}}}\leqslant 6\sqrt{\frac{{{u_{n}}(\delta)}}{{{p_{k+m+1}}}}}.$
B.2 Proof of Theorem B.1
The following Lemma B.2 shows that if $k_{n}-m_{n}\to+\infty$, then $\beta\left({{k_{n}},{m_{n}}}\right)\to\beta$. Theorem B.1 now follows by the same argument used to prove Theorem 2.1, using the convergence of $\beta\left({{k_{n}},{m_{n}}}\right)$ instead of Lemma A.3 and Lemma B.1 instead of Lemma A.1.
Lemma B.2.
Let $\alpha_{n}$, $k_{n}$ and $b_{n}$ be sequences such that, $\alpha_{n}\to\alpha$, $k_{n}\to+\infty$ and $k_{n}-n_{n}\to+\infty$. Then,
$\frac{1}{{2{b_{n}}+1}}\sum\limits_{j=-{b_{n}}}^{{b_{n}}}{{\alpha_{{k_{n}}+j}}}\to\alpha.$
Proof.
Let ${\rho_{n}}=\frac{1}{{2{b_{n}}+1}}\sum\limits_{j=-{b_{n}}}^{{b_{n}}}{{\alpha_{{k_{n}}+j}}}$, then
${\rho_{n}}-\alpha=\frac{1}{{2{b_{n}}+1}}\sum\limits_{j=-{b_{n}}}^{{b_{n}}}{\left({{\alpha_{{k_{n}}+j}}-\alpha}\right)}=\frac{1}{{2{b_{n}}+1}}\sum\limits_{j=0}^{2{b_{n}}}{\left({{\alpha_{{k_{n}}-{b_{n}}+j}}-\alpha}\right)}.$
Fix $\epsilon>0$. The convergence of $\alpha_{n}$ ensures the existence of $N_{1}$ such that $\left|{{\alpha_{n}}-\alpha}\right|<\epsilon$ for all $n\geq{N_{1}}$. Given that $k_{n}-b_{n}\to+\infty$, there exists $N_{2}$ satisfying ${k_{n}}-{b_{n}}\geq{N_{1}},$ for all $n\geq N_{2}$, which yields $\left|{{\alpha_{{k_{n}}-{b_{n}}+j}}-\alpha}\right|\leq\epsilon$ for all $n\geq N_{2}$ and $j\in\mathbb{N}$. Consequently, for $n\geq N_{2}$,
$\left|{{\rho_{n}}-\alpha}\right|\leq\frac{1}{{2{b_{n}}+1}}\sum\limits_{j=0}^{2{b_{n}}}{\left|{{\alpha_{{k_{n}}-{b_{n}}+j}}-\alpha}\right|}\leq\frac{1}{{2{b_{n}}+1}}\sum\limits_{j=0}^{2{b_{n}}}\epsilon=\epsilon.$
B.3 Proof of Corollary B.1
We just need to show that sequences $k_{n}=A\ln n$ and $m_{n}={A\ln n}/{l}$ satisfy conditions (i), (ii) and (iii) of Theorem B.1. The first two are trivially satisfied, for the third one, notice that
$\mathop{\lim}\limits_{n}\frac{{{e^{\left({A\ln n+\frac{{A\ln n}}{l}}\right)\beta}}}}{{L\left({{e^{A\ln n+\frac{{A\ln n}}{l}}}}\right)}}\frac{{\ln n}}{n}=\mathop{\lim}\limits_{n}\frac{1}{{{n^{\frac{{1-\left({1+\frac{1}{l}}\right)A\beta}}{2}}}L\left({{n^{\left({1+\frac{1}{l}}\right)A}}}\right)}}\frac{{\ln n}}{{{n^{\frac{{1-\left({1+\frac{1}{l}}\right)A\beta}}{2}}}}}.$
The condition $l>{A\beta}/{(1-A\beta)}$ implies that $01-\left({1+{1}/{l}}\right)A\beta>$ 0, therefore,
$\mathop{\lim}\limits_{n}\frac{1}{{{n^{\frac{{1-\left({1+\frac{1}{l}}\right)A\beta}}{2}}}L\left({{n^{\left({1+\frac{1}{l}}\right)}}}\right)}}=0\quad\textnormal{and}\quad\mathop{\lim}\limits_{n}\frac{{\ln n}}{{{n^{\frac{{1-\left({1+\frac{1}{l}}\right)A\beta}}{2}}}}}=0,$
which shows that $k_{n}$ and $m_{n}$ satisfy condition (iii) in Theorem B.1.
Appendix C Tail index estimation in the positive recurrent case
In this section, we prove the asymptotic normality of our estimator in the positive recurrent case. The main result is the following
Theorem C.1 (Asymptotic normality in the positive recurrent case).
Suppose $X$ is a regenerative positive recurrent Markov chain such that the distribution of its regeneration time satisfies equation (1). Assume that $k_{n}$ satisfies the hypothesis of Theorem 2.2 as well as the following extra assumption:
  • (A4)
    There exists $D^{\prime}>0$ such that for any $0<D<D^{\prime}$
    $\sum_{j=n+1}^{n\left(1+D\right)}{\bar{F}\left(e^{k_{j}}\right)}=nD\bar{F}\left(e^{k_{n}}\right)+o\left(\bar{F}\left(e^{k_{n}}\right)\right).$ (46)
Then,
  • (i)
    as $n\to+\infty$, we have the convergence in distribution:
    $\sqrt{{N_{n}}p_{k_{N_{n}}}}\left(\widehat{\beta}_{N_{n}}({k_{N_{n}}})-\beta(k_{N_{n}})\right)\Rightarrow\mathcal{N}\left(0,\;e^{\beta}-1\right).$
  • (ii)
    In addition, asymptotic normality holds true for the ’standardized’ deviation:
    $\frac{{\sqrt{{N_{n}}\widehat{p}_{{k_{N_{n}}}}^{({N_{n}})}}\left({{{\widehat{\beta}}_{N_{n}}}\left({{k_{N_{n}}}}\right)-\beta\left({{k_{N_{n}}}}\right)}\right)}}{{\sqrt{{e^{{{\widehat{\beta}}_{N_{n}}}\left({{k_{N_{n}}}}\right)}}-1}}}\Rightarrow\mathcal{N}\left({0,\;1}\right),\text{ as }n\to+\infty.$
Remark 7.
Assumption (A4) can be replaced by the following slightly more restrictive, but easier to verify, assumption:
  • (A4’)
    There exists $D^{\prime}>0$ such that for any $0<D<D^{\prime}$
    $\bar{F}\left(e^{k_{n(1+D)}}\right)=\bar{F}\left(e^{k_{n}}\right)+o\left(\frac{\bar{F}\left(e^{k_{n}}\right)}{n}\right).$
As in the case of Theorem 2.2, the main result follows directly from the following lemma, which is an extension of Lemma A.7 for the case where the number of i.i.d samples we have is random.
Lemma C.1.
Let $W_{n}$ be a sequence of i.i.d. random variables with survival function (1), and $k_{n}$ be a sequence that satisfies the hypothesis of Theorem C.1. Suppose that $T_{n}$ is a sequence of positive, integer-valued random variables such that $\frac{T_{n}}{n}$ converges in probability to some positive number $\theta$. Then
$\sqrt{T_{n}\bar{F}\left({{e^{{k_{T_{n}}}}}}\right)}\left({\frac{{\sum\limits_{i=1}^{T_{n}}{\mathbb{I}\left\{{{W_{i}}>{e^{{k_{T_{n}}}}}}\right\}}}}{{\sum\limits_{i=1}^{T_{n}}{\mathbb{I}\left\{{{W_{i}}>{e^{{k_{T_{n}}}+1}}}\right\}}}}-\frac{{\bar{F}\left({{e^{{k_{T_{n}}}}}}\right)}}{{\bar{F}\left({{e^{{k_{T_{n}}}+1}}}\right)}}}\right).$
converges weakly to a centred normal distribution with variance ${e^{2\beta}}\left({{e^{\beta}}-1}\right)$.
C.1 Proof of Lemma C.1
For this proof, we will reuse the notation we utilized in the proof of Lemma A.7 and will add the following definitions: $q_{n}=nv_{n}$, ${y_{n}}=\sqrt{{{{v_{n}}}}/{{(n{w_{n}}^{2})}}}$ and
$U_{n,k}$ $=\sum\limits_{i=1}^{n}{\left({{\phi_{k}}\left({{W_{n,i}}}\right)-\mathbb{E}\left[{\phi_{k}}\left({{W_{n,i}}}\right)\right]}\right)}.$
By equation (31) and Lemma A.6
$\sqrt{q_{n}}\left({\widehat{\lambda}_{n}}-{\lambda_{n}}\right)={y_{n}}\sqrt{q_{n}}\left({{{\widetilde{Z}}_{n}}\left({{\phi_{1}}}\right)-{\lambda_{n}}{{\widetilde{Z}}_{n}}\left({{\phi_{2}}}\right)}\right)+{o_{P}}\left(1\right),$
and ${({{{\widetilde{Z}}_{n}}\left({{\phi_{k}}}\right)})_{1\leqslant k\leqslant 2}}$ converges weakly to a centred normal distribution. Using that $\lambda_{n}\to e^{\beta}$ and $y_{n}\sim{e^{\beta}}/{\sqrt{q_{n}}}$, this implies that
$\frac{\sqrt{q_{n}}\left({\widehat{\lambda}_{n}}-{\lambda_{n}}\right)}{e^{\beta}}={{{\widetilde{Z}}_{n}}\left({{\phi_{1}}}\right)-e^{\beta}{{\widetilde{Z}}_{n}}\left({{\phi_{2}}}\right)}+{o_{P}}\left(1\right).$ (47)
Take $n_{0}=\left\lfloor\theta n\right\rfloor$, and $V_{n}=\sqrt{q_{n}}({{{\widetilde{Z}}_{n}}\left({{\phi_{1}}}\right)-e^{\beta}{{\widetilde{Z}}_{n}}\left({{\phi_{2}}}\right)})=U_{n,1}-e^{\beta}U_{n,2}$, then
$\frac{\sqrt{q_{T_{n}}}}{e^{\beta}}\left({\widehat{\lambda}_{T_{n}}}-{\lambda_{T_{n}}}\right)=\left(\frac{V_{n_{0}}}{\sqrt{q_{n_{0}}}}+\frac{V_{T_{n}}-V_{n_{0}}}{\sqrt{q_{n_{0}}}}\right)\sqrt{\frac{q_{n_{0}}}{q_{T_{n}}}}+o_{P}(1).$
By Lemma A.7 and our assumption about the convergence in probability of ${T_{n}}/{n}$, we have that ${V_{n_{0}}}/{\sqrt{q_{n_{0}}}}$ converges in distribution to a centred Normal random variable with variance $e^{2\beta}({e^{\beta}}-1)$ and ${q_{n_{0}}}/{q_{T_{n}}}$ converges in probability to $1$. Therefore, if we show that
$\frac{V_{T_{n}}-V_{n_{0}}}{\sqrt{q_{n_{0}}}}\mathop{\to}0$ (48)
in probability then our lemma will be proved by two successive applications of Slutsky’s theorem.
Notice that $V_{T_{n}}-V_{n_{0}}=U_{T_{n},1}-U_{n_{0},1}-e^{\beta}\left(U_{T_{n},2}-U_{n_{0},2}\right)$, hence, if we show that ${(U_{T_{n},1}-U_{n_{0},1})}/{\sqrt{q_{n_{0}}}}$ and ${(U_{T_{n},2}-U_{n_{0},2})}/{\sqrt{q_{n_{0}}}}$ converge to 0 in probability, then (48) will be proved. Given that the proofs of both convergences are analogous, we will only demonstrate the first one.
Let $\epsilon>0$ be fixed, and set $n_{1}=\left\lfloor n_{0}\left(1-{\epsilon^{3}}/{32}\right)\right\rfloor+1,n_{2}=\left\lfloor n_{0}\left(1+{\epsilon^{3}}/{32}\right)\right\rfloor$, then
$\mathbb{P}\left(\left|U_{T_{n},1}-U_{n_{0},1}\right|>\epsilon\sqrt{q_{n_{0}}}\right)\leq I_{n,1}+I_{n,2},$ (49)
where,
$I_{n,1}$ $=\mathbb{P}\left(T_{n}\notin[n_{1},n_{2}]\right),$
$I_{n,2}$ =($=\mathbb{P}\left(\left\{\left|U_{T_{n},1}-U_{n_{0},1}\right|>\epsilon\sqrt{q_{n_{0}}}\right\}\cap T_{n}\in[n_{1},n_{2}]\right).$
The convergence in probability of ${T_{n}}/{n}$ to $\theta$ implies that there exists $N_{1}$ such that $I_{n,1}<{\epsilon}/{2}$ for all $n\geq N_{1}$, hence,
$\forall n\geq N_{1}\quad\mathbb{P}\left(\left|U_{T_{n},1}-U_{n_{0},1}\right|>\epsilon\sqrt{q_{n_{0}}}\right)\leq\frac{\epsilon}{2}+I_{n,2}.$ (50)
To bound the second term on the right-hand side of the previous display, observe that $I_{n,2}$ is smaller than
$\mathbb{P}\left(\max_{n_{1}\leq j\leq n_{0}}\left|U_{j,1}-U_{n_{0},1}\right|>\epsilon\sqrt{q_{n_{0}}}\right)+\mathbb{P}\left(\max_{n_{0}<j\leq n_{2}}\left|U_{j,1}-U_{n_{0},1}\right|>\epsilon\sqrt{q_{n_{0}}}\right).$
We just need to focus on the case $n_{0}<j\leq n_{2}$ because the other will be analogous. To ease the notation, for any $a<b$, we will write $\bar{F}_{a}^{b}$ instead of $\bar{F}\left(u_{a}\right)-\bar{F}\left(u_{b}\right)$ and we will use $\bar{F}_{a}$ to denote $\bar{F}(u_{a})$. Let $\mathcal{C}_{n}$ be the set $\{\max_{n_{0}<j\leq n_{2}}|U_{j,1}-U_{n_{0},1}|>\epsilon\sqrt{q_{n_{0}}}\}$. We can write the difference $U_{j,1}-U_{n_{0},1}$ as
$U_{j,1}-U_{n_{0},1}$ $=\sum\limits_{i=1}^{j}\left(\mathbb{I}\left\{{{W_{i}}>u_{j}}\right\}-\bar{F}_{j}\right)-\sum\limits_{i=1}^{n_{0}}\left(\mathbb{I}\left\{{{W_{i}}>u_{n_{0}}}\right\}-\bar{F}_{n_{0}}\right)$
$=\sum\limits_{i=n_{0}+1}^{j}\left(\mathbb{I}\left\{{{W_{i}}>u_{n_{0}}}\right\}-\bar{F}_{n_{0}}\right)-\sum\limits_{i=1}^{j}\left(\mathbb{I}\left\{{{W_{i}}\in(u_{n_{0}},u_{j}]}\right\}-\bar{F}_{n_{0}}^{j}\right).$ (51)
Suppose for the moment that
$\frac{\underset{n_{0}<j\leq n_{2}}{\max}\left|\sum\limits_{i=1}^{j}\left(\mathbb{I}\left\{{{W_{i}}\in(u_{n_{0}},u_{j}]}\right\}-\bar{F}_{n_{0}}^{j}\right)\right|}{\sqrt{q_{n_{0}}}}\to 0$ (52)
in probability.
Let $\mathcal{G}_{n}$ be the event $\{\max_{n_{0}<j\leq n_{2}}{|\sum_{i=1}^{j}(\mathbb{I}\left\{{{W_{i}}\in(u_{n_{0}},u_{j}]}\right\}-\bar{F}_{n_{0}}^{j})|}\leq{\epsilon\sqrt{q_{n_{0}}}}/{2}\}$. By (52), we can find $N_{2}$ such that $\mathbb{P}\left(\mathcal{G}_{n}\right)\geq 1-{\epsilon}/{8}$ for all $n\geq N_{2}$. Therefore, for all $n\geq N_{2}$ we have
$\mathbb{P}\left(\mathcal{C}_{n}\right)\leq\mathbb{P}\left(\left\{\max_{n_{0}<j\leq n_{2}}\left|U_{j,1}-U_{n_{0},1}\right|>\epsilon\sqrt{q_{n_{0}}}\right\}\cap\mathcal{G}_{n}\right)+\frac{\epsilon}{8}.$
Using (51), we obtain that $\{\max_{n_{0}<j\leq n_{2}}\left|U_{j,1}-U_{n_{0},1}\right|>\epsilon\sqrt{q_{n_{0}}}\}\cap\mathcal{G}_{n}$ is contained in the event
$\left\{\underset{n_{0}<j\leq n_{2}}{\max}\left|\sum\limits_{i=n_{0}+1}^{j}\left(\mathbb{I}\left\{{{W_{i}}>u_{n_{0}}}\right\}-\bar{F}_{n_{0}}\right)\right|>\frac{\epsilon\sqrt{q_{n_{0}}}}{2}\right\}.$
By Kolmogorov inequality,
$\mathbb{P}\left(\underset{n_{0}<j\leq n_{2}}{\max}\left|\sum\limits_{i=n_{0}+1}^{j}\left(\mathbb{I}\left\{{{W_{i}}>u_{n_{0}}}\right\}-\bar{F}_{n_{0}}\right)\right|>\frac{\epsilon\sqrt{q_{n_{0}}}}{2}\right)\leq\frac{4\left(n_{2}-n_{0}\right)\bar{F}_{n_{0}}}{\epsilon^{2}n_{0}\bar{F}_{n_{0}}}.$
The right-hand side of the previous equation equals ${4\left(\left\lfloor n_{0}\left(1+{\epsilon^{3}}/{32}\right)\right\rfloor-n_{0}\right)}/{(\epsilon^{2}n_{0})}$, and that is smaller than $\epsilon/8$. Hence, $\mathbb{P}\left(\mathcal{C}_{n}\right)\leq\frac{\epsilon}{4}$ for all $n\geq N_{2}$. In a similar fashion, we can find $N_{3}$ such that $\mathbb{P}\left(\max_{{n_{1}\leq j\leq n_{0}}}\left|U_{j,1}-U_{n_{0},1}\right|>\epsilon\sqrt{q_{n_{0}}}\right)\leq{\epsilon}/{4}$ for $n\geq N_{3}$. This shows that $I_{n,2}\leq{\epsilon}/{2}$ for $n\geq\max(N_{2},N_{3})$. Combining this with equation (50), proofs (48).
To finish, we proceed with the proof of (52). Let $\delta>0$ be fixed. Without loss of generality, assume that $21+{\epsilon^{3}}/{8}<$ 2 and ${\epsilon^{3}}/{8}<D^{\prime}$. Denote by $\mathcal{H}_{n,\delta}$ the event $\{{\max_{n_{0}<j\leq n_{2}}}|\sum_{i=1}^{j}(\mathbb{I}\{{{W_{i}}\in(u_{n_{0}},u_{j}]}\}-\bar{F}_{n_{0}}^{j})|>\delta\sqrt{q_{n_{0}}}\}$, then,
$\mathcal{H}_{n,\delta}\subseteq\bigcup_{j=n_{0}+1}^{n_{2}}{\left\{\left|\sum\limits_{i=1}^{j}\left(\mathbb{I}\left\{{{W_{i}}\in(u_{n_{0}},u_{j}]}\right\}-\bar{F}_{n_{0}}^{j}\right)\right|>\delta\sqrt{q_{n_{0}}}\right\}}.$ (53)
By Chebyshev’s inequality,
$\mathbb{P}\left(\left|\sum\limits_{i=1}^{j}\left(\mathbb{I}\left\{{{W_{i}}\in(u_{n_{0}},u_{j}]}\right\}-\bar{F}_{n_{0}}^{j}\right)\right|>\delta\sqrt{q_{n_{0}}}\right)\leq\frac{2\bar{F}_{n_{0}}^{j}}{\delta^{2}\bar{F}_{n_{0}}}=\frac{2}{\delta^{2}}\left(1-\frac{\bar{F}_{j}}{\bar{F}_{n_{0}}}\right).$
Combining this with (53) and (A4), we obtain
$\mathbb{P}\left(\mathcal{H}_{n,\delta}\right)$ $\leq\frac{2}{\delta^{2}}\sum_{j=n_{0}+1}^{n_{2}}{\left(1-\frac{\bar{F}_{j}}{\bar{F}_{n_{0}}}\right)}\leq\frac{2}{\delta^{2}}\left(n_{2}-n_{0}-\frac{\sum_{j=n_{0}+1}^{n_{2}}\bar{F}_{j}}{\bar{F}_{n_{0}}}\right)$
$\leq\frac{2}{\delta^{2}}\left(n_{2}-n_{0}-\frac{\left(n_{2}-n_{0}\right){\bar{F}_{n_{0}}+o\left({\bar{F}_{n_{0}}}\right)}}{{\bar{F}_{{n_{0}}}}}\right)=o\left(1\right),$
which completes the proof of (52).
Appendix D Functional version of the limit distributions
Let $n\geq 1$, and define the step function
$\sigma_{n}(B):\;t\geq 0\mapsto\frac{\Sigma_{\lfloor{nt}\rfloor}(B)}{n^{\beta}L_{v,D}(n)}.$ (54)
The result stated below describes the asymptotic distribution of the process $\sigma_{n}(B)$. See Theorem 17.4.4 in [31] for an analogous result in the positive recurrent case. We denote by $M_{\beta}=(M_{\beta}(t))_{t\geq 0}$ the Mittag-Leffler process with parameter $\beta\in(0,1)$, defined by:
$\mathbb{E}\left[\left(M_{{}_{\beta}}(1)\right)^{m}\right]$ $=\frac{{m!}}{{\Gamma\left({1+m\beta}\right)}},\quad\text{for all }m\geqslant 0,$
${M_{\beta}}\left(t\right)$ $\mathop{=}\limits^{d}{t^{\beta}}{M_{\beta}}\left(1\right),\quad\text{for all }t\geq 0.$
The characteristic functions describing the marginal distributions are given by (see 3.39 in [25])
$\mathbb{E}\left[{{e^{i\zeta{M_{\beta}}\left(t\right)}}}\right]=\sum\limits_{k=0}^{+\infty}{\frac{{{{\left({i\zeta{t^{\beta}}}\right)}^{k}}}}{{\Gamma\left({1+k\beta}\right)}}},\quad\zeta\in\mathbb{R},\;t\geqslant 0.$ (55)
Theorem D.1 (Functional Limit Theorem).
Let $\beta\in(0,1)$ and $\nu$ be any probability distribution. Suppose that the chain $X$ is $\beta$-regular and $B$ is a Harris set. Then, as $n\to\infty$, we have:
[…]
in the sense of Skorokhod topology.
Remark 8.
Let $Y=1/Z_{\beta}^{\beta}$ where $Z_{\beta}$ is as in Theorem A.1. By virtue of equation (8.3) in page 453 of [14], the Laplace transform of $Y$ is
$\mathbb{E}\left[{{e^{-sY}}}\right]=\sum\limits_{k=0}^{+\infty}{\frac{{{{\left({-\Gamma\left({1+\beta}\right)s}\right)}^{k}}}}{{\Gamma\left({1+k\beta}\right)}}},$
which equals the Laplace transform of $\Gamma\left({1+\beta}\right){M_{\beta}}\left(1\right)$,  cf (55).
D.1 Proof of Theorem D.1
Let $L_{s}\left(n\right)$ be such that $L\left(n\right)={(\Gamma\left({1-\beta}\right){L_{s}}\left(n\right))}^{-1}$ with $L$ as defined in (1). Observe that,
${G_{\nu,D}}\left(n\right)=\frac{1}{{\mu\left(D\right)}}\sum\limits_{k=1}^{n}{{\mathbb{P}_{\nu}}\left({{X_{k}}\in D}\right)}=\frac{1}{{\mu\left(D\right)}}{\mathbb{E}_{\nu}}\left[{\sum\limits_{k=1}^{n}{\mathbb{I}\left\{{{X_{k}}\in D}\right\}}}\right].$
By Lemma 3.1 and Definition 3.2 of [25] ${\mathbb{E}_{\nu}}\left[{\sum_{k=1}^{n}{\mathbb{I}\left\{{{X_{k}}\in D}\right\}}}\right]$ is asymptotically equivalent to $\Gamma\left({1+\beta}\right)^{-1}{n^{\beta}}\mu\left(D\right){L_{s}}\left(n\right),$ therefore,
${G_{\nu,D}}\left(n\right)\sim\frac{1}{{\Gamma\left({1+\beta}\right)}}{n^{\beta}}{L_{s}}\left(n\right).$ (56)
Let $u\left(n\right)=n^{\beta}L_{s}\left(n\right)$, and define the process
${{\mathbf{T}}_{n,B}}={\left\{{\frac{{{S_{\left\lfloor{nt}\right\rfloor}}\left(B\right)}}{{u\left(n\right)}}}\right\}_{t\geqslant 0}}=\frac{{{G_{\nu,D}}\left(n\right)\mu\left(B\right)}}{{u\left(n\right)}}{{\mathbf{S}}_{n,B}}.$
By Lemma 3.6 and Theorem 3.2 of [25], ${{\mathbf{T}}_{n,B}}$ converges weakly on the Skorokhod topology to the process $\mu\left(B\right)M_{\beta}$ and by (56), ${{{G_{\nu,D}}\left(n\right)}}/{{u\left(n\right)}}\to{1}/{{\Gamma\left({1+\beta}\right)}}$ which completes the proof.
Appendix E Nummelin’s splitting technique and construction of pseudo regeneration blocks
Nummelins splitting technique relies crucially on the notion of small set: a set $K\in\mathcal{E}$ is said to be small if there exist $m\in\mathbb{N}^{\ast}$, $\delta>0$ and a probability measure $\Phi$ supported by $K$ such that
$\forall\left(x,B\right)\in K\times\mathbb{E},\;\;\Pi_{m}\left(x,B\right)\geq\delta\Phi\left(B\right).$ (57)
[…]
  • for any $x\notin K$, $B\in\mathcal{E}$, $b$ and $b^{\prime}$ in $\left\{0,1\right\},$
    $\Pi_{\mathcal{M}}\left(\left(x,b\right),B\times\left\{b^{\prime}\right\}\right)=\Pi\left(x,B\right)\times\text{Ber}_{\delta}(b^{\prime}),$
  • for any $x\in K$, $B\in\mathcal{E}$, $b^{\prime}$ in $\left\{0,1\right\}$,
    $\left\{\begin{array}[]{ccl}\Pi_{\mathcal{M}}\left(\left(x,1\right),B\times\left\{b^{\prime}\right\}\right)&=&\Phi(B)\times\text{Ber}_{\delta}(b^{\prime}),\\ \Pi_{\mathcal{M}}\left(\left(x,0\right),B\times\left\{\beta^{\prime}\right\}\right)&=&(1-\delta)^{-1}(\Pi\left(x,B\right)-\delta\Phi(B))\times\text{Ber}_{\delta}(b^{\prime}).\end{array}\right.$
The key point of the construction relies on the fact that $A_{K}=K\times\{1\}$ is an atom for the bivariate chain $X^{\mathcal{M}}=(X,Y)$, which inherits all its communication and stochastic stability properties from $X$.
Recall also that conditions of type (57) can be replaced by Foster-Lyapunov drift conditions that are much more tractable in practice, see e.g. Chapter 11 in [31]. The construction above permits the extension of probabilistic results established for regenerative chains to general recurrent Harris chains. In particular, we have the following result, presented in [8, pp 19].
Proposition E.1.
Let $X$ be a $\beta$-regular chain, with $\beta\in[0,1]$. Suppose that condition $\mathcal{M}=\mathcal{M}(1,K,\delta,\Psi)$ is fulfilled, then, the split chain $X^{\mathcal{M}}$ is $\beta$-regular.
Hence, $X$’s regularity index $\beta$ is the regular variation index of the conditional survivor function of the hitting time $\tau_{A_{K}}=\inf\{n\geq 1:\;(X_{n},Y_{n})\in A_{K}\}$ given $(X_{0},Y_{0})\in A_{K}$. However, the return times to $A_{K}$ are not observable, just like the sample path of the Nummelin extension, and cannot be straightforwardly exploited from a statistical perspective. We shall explain in subsection 3.5 how estimators tailored to the regenerative case can be nevertheless extended to the pseudo-regenerative case in practice by means of the plug-in approximation procedure originally proposed in [6] in the positive recurrent case.
E.1 Obtaining samples form the split chain
Proposition E.1 and the algorithmic construction described after Eq. (57) guarantee that if the chain satisfies the minorization condition $\mathcal{M}=\mathcal{M}(1,K,\delta,\Psi)$, and $K,\delta$ and $\Psi$ are known, then we can generate samples of the split chain, which is atomic and has the same $\beta$ as the original chain. Assume the existence of a $\sigma$-finite measure $\lambda$ of reference on $(E,\mathcal{E})$ that dominates the conditional probability measures $\Pi(x,\;dy)$, $x\in E$, and the initial distribution $\nu$: $\Pi(.,dy)=\pi(.,y)\lambda(dy)$ and $\nu(dy)=g(y)\lambda(dy)$. Notice incidentally that the measure $\Psi$ involved in $\mathcal{M}$ is then absolutely continuous w.r.t. $\lambda$ as well: $\Psi(dy)=\psi(y)\lambda(dy)$ and then $\pi\left(x,y\right)\geq\delta\psi\left(y\right)$ for all $(x,y)\in K^{2}$. As shown in Section 3.2 in [6], given $X^{(n+1)}=\left(X_{1},\ldots,X_{n+1}\right)$, samples from the distribution of $Y^{(n)}=\left(Y_{1},\ldots\right.$, $\left.Y_{n}\right)$ can be obtained as follows. From $i=1$ to $n$, the r.v. $Y_{i}$ is drawn from a Bernoulli distribution with parameter $\delta$, unless $X$ hits the small set $K$ at time $i$: in the latter case, $Y_{i}$ is drawn from a Bernoulli distribution with parameter $\delta\psi\left(X_{i+1}\right)/\pi\left(X_{i},X_{i+1}\right)$. Given that $A_{K}=K\times\{1\}$ is an atom for the split chain, and the statistics under study in this paper only depends on the size of the regeneration blocks, sampling $Y_{i}$ when $X_{i}\in K$ is sufficient here. The accuracy of the estimator improves as the (random) number of samples (the number of regeneration blocks namely) increases. This number is influenced by the size of the chosen small set and how frequently the chain visits it in a finite-length trajectory. It is also affected by the sharpness of the lower bound in the minorization condition. Essentially, there is a trade-off that can be described as follows. Increasing the size of the small set $K$ used for constructing the pseudo-blocks naturally increases the number of time points that could determine a block (or a cut in the trajectory). However, it also reduces the probability of cutting the trajectory, as the uniform lower bound for $\pi(x,y)$ over $K^{2}$ then decreases. This suggests a criterion for selecting the small set $K$: choose a small set that maximizes the maximum expected number of data blocks given the trajectory, that is
$N_{n}(K)=\mathbb{E}_{\nu}\left[\sum_{i=1}^{n}\mathbb{I}\left\{X_{i}\in K,Y_{i}=1\right\}\mid X^{(n+1)}\right].$
In Section 3.6 of [6], a data-driven approach to select the small set is proposed for the cases where the chain takes real values. The idea relies on the fact that, in many cases, for a well-chosen $x_{0}$ and $\epsilon$ small enough, certain intervals $V_{x_{0},\epsilon}=[x_{0}-\epsilon,x_{0}+\epsilon]$ are small sets, with the minorization measure $\Psi$ being the Lebesgue measure on $V_{x_{0},\epsilon}$. Given a point $x_{0}$ (generally taken as the mean or the median of the $X_{i}$’s), the proposed algorithm finds the value of $\epsilon$ that maximizes the expected number of regeneration blocks, that is
$N_{n}\left(V_{x_{0},\epsilon}\right)=\frac{\delta\left(V_{x_{0},\epsilon}\right)}{2\epsilon}\sum_{i=1}^{n}\frac{\mathbb{I}\left\{\left(X_{i},X_{i+1}\right)\in V_{x_{0},\epsilon}^{2}\right\}}{\pi\left(X_{i},X_{i+1}\right)},$
where $\delta\left(V_{x_{0},\epsilon}\right)=2\epsilon\inf_{(x,y)\in V^{2}_{x_{0},\epsilon}}\pi(x,y)$. Then, the samples of the split chain can be obtained by following the procedure described at the begining of this subsection with $K=V_{x_{0},\epsilon}$, $\delta=2\epsilon\inf_{(x,y)\in V^{2}_{x_{0},\epsilon}}\pi(x,y)$ and $\psi(y)=1/(2\epsilon)$.